### v2

# Univalent Material Set Theory

Abstract
Homotopy type theory (HoTT) can be seen as a generalisation of structural set theory, in the sense that 0-types represent structural sets within the more general notion of types. For material set theory, we also have concrete models as 0-types in HoTT, but this does not currently have any generalisation to higher types. The aim of this paper is to give such a generalisation of material set theory to higher type levels within homotopy type theory. This is achieved by generalising the construction of the type of iterative sets [gylterud-iterative] to obtain an $n$-type universe of $n$-types. At level 1, this gives a connection between groupoids and multisets.
More specifically, we define the notion of an $\in$-structure as a type with an extensional binary type family and generalise the axioms of constructive set theory to higher type levels. There is a tight connection between the univalence axiom and the extensionality axiom of $\in$-structures. Once an $\in$-structure is given, its elements can be seen as representing types in the ambient type theory. A useful property of these structures is that an $\in$-structure of n-types is itself an n-type, as opposed to univalent universes, which have higher type levels than the types in the universe.
The theory has an alternative, coalgebraic formulation, in terms of coalgebras for a certain hierarchy of functors, $P^{n}$, which generalises the powerset functor from sub-types to covering spaces and $n$-connected maps in general. The coalgebras which furthermore are fixed-points of their respective functors in the hierarchy are shown to model the axioms given in the first part.
As concrete examples of models for the theory developed we construct the initial algebras of the $P^{n}$ functors. In addition to being an example of initial algebras of non-polynomial functors, this construction allows one to start with a univalent universe and get a hierarchy of $\in$-structures which gives a stratified $\in$-structure representation of that universe. These types are moreover $n$-type universes of $n$-types which contain all the usual types an type formers. The universes are cumulative both with respect to universe levels and with respect to type levels.
All the results are formalised in the proof-assistant Agda.
1 Introduction
Material set theories are set theories which emphasise the notion of sets as collections of elements (often themselves sets) and where the identity of individual elements is tracked across sets, usually with a global, binary membership relation ($\in$). This category includes the traditional Zermelo–Fraenkel (ZF) set theory. ZF is a theory in the language of first-order logic, intended as a foundation for mathematics. In what follows, we work within the framework of homotopy type theory (HoTT), fully formalized in Agda. HoTT is a structural framework, with the Univalence Axiom in particular allowing identification of types which are structurally the same, i.e. equivalent – deemphasising the individual elements and their identity outside the structure of the type. Taking a step back, one can see HoTT as a generalisation of structural set theory to higher type levels. The high-level question we attempt to answer in this paper is: What is the corresponding generalisation of material set theory to higher type levels? From this vantage point we will regard material set theory in the same way that a classical mathematician regards group theory. Namely, we study a certain type of mathematical structures, and interest ourselves in their properties and relationships. The structures we study are not groups, but what we call $\in$-structures: types with an extensional elementhood relation.
Since we are working in HoTT, we can consider the type level of the underlying type of sets and of the relation of an $\in$-structure. In classical set theory, the statement “$x\in y$” is a proposition. But in our setting, we can consider $\in$-structures where $x\in y$ is a type of any type level. Similarly, we can generalise the axioms of ZF to higher type levels. An example of such a $\in$-structure, where $x\in y$ is allowed to be a type of any level, was considered in Multisets in type theory [gylterud-multisets] by one of the authors.
We aim to give a higher level generalisation of material set theory, by considering $\in$-structures where $x\in y$ is an $n$-type. With care, the usual properties, which we know and love from set-theory, can be reformulated and proven to hold in our models. But sometimes what used to be a single property generalises to several properties when taking higher type levels into consideration. Let us, for the sake of building some intuition, take a closer look at level 1 of this generalisation.
In a level 1 $\in$-structure, elements are related by a set-valued $\in$-relation: given two elements, $x,y:V$ the type $x\in y$ is a set. One way of understanding this to think of $y$ as a multiset where $x$ may occur more than once. For instance, if $x\in y$ is a finite type with $n$ elements, then we can think of this as saying that $x$ occurs $n$ times in $y$. The generalised properties support this interpretation: For instance, level 1 (unordered) tupling allows the formation of multisets of the form $\{x_{0},\cdots,x_{n}\}_{1}$ where repetitions will be counted separately. But a level 1 $\in$-structure may also support level 0 (unordered) tupling, with a separate operation, $\{x_{0},\cdots,x_{n}\}_{0}$, which becomes a set: $x\in\{x_{0},\cdots,x_{n}\}_{0}$ being proposition for any $x$.
[…]
There is an immediate connection between univalent set theory and homotopy type theory, whereby there is an equivalence between $\in$-structures and coalgebras for the $n$-truncated maps functors $\operatorname{P}^{n+1}_{U}:\operatorname{Type}\rightarrow\operatorname{Type}$, which associates to each type $X$ the type of n-truncated maps into $X$. Thus, $\operatorname{P}^{0}_{U}X$ is the type of subtypes of $X$, while $\operatorname{P}^{1}_{U}X$ is the type of covering spaces of $X$ and so on. We show that these functors have initial algebras, $\operatorname{V}^{n}$, and determine the univalent set theory axioms satisfied by the initial algebras and other fixed-points of these functors. These initial algebras hence form a family of models of univalent material set theory, motivating the axioms and interpolating between the standard iterative hierarchy and the generalised multisets.
One way in which univalent material set theory distinguishes itself in HoTT is that type levels are off by one. What this means is that models form $n$-type based families of $n$-types: if $A,B:V$ are $n$-types in $(V,\in)$ (a notion made precise in Definition 4) then $A=B$ is an $n-1$-type. This means that $V$ is an $n$-type. This contrasts the situation for univalent universes, where a well-known result [kraus_truncation_2015] states that if $U$ contains strict $n$-types the type level of $U$ is itself at least $n+1$.
The model given by the initial algebra $\operatorname{V}^{n}$ is of level $n$, and thus the type $\operatorname{V}^{n}$ is an $n$-type. The type can be equipped with the structure of a Tarski style universe. The decoding of an element in $\operatorname{V}^{n}$ is an $n$-type, so $\operatorname{V}^{n}$ is an $n$-type universe of $n$-types. Moreover, the decoding holds up to definitional equality, making it very ergonomic to use.
Type levels being off by one might seem strange at first, but it is caused by the $\in$-relation imposing extra structure and thereby killing automorphisms. This observation generalises what is known about the cumulative hierarchy in models of (usual) set theory in HoTT, where $V$ is a set of sets. Especially in category theory, this can be useful to strictify structures. For instance, as explored in The Category of Iterative Sets in Homotopy Type Theory and Univalent Foundations [gratzer2024category], when recreating the category with family structure on sets in HoTT, one finds oneself blocked by the fact that the types in a context forms a strict groupoid, not a set. By using an $\in$-structure as the category of contexts, the off-by-one property sidesteps this block, yielding a good category with families.
Another perspective we explore is extracting types from $\in$-structures. A given element of an $\in$-structure has a type of elements, and considering the whole $\in$-structure we can ask what types can be represented as types of elements within it. Some insights into how replacement affects representations of types, such as $\mathbb{N}$ for the axiom of infinity, has been collected in Section 3. In particular, we show that the replacement property in set theory says that the $\in$-structure supports all choices of representations of a type equally (Proposition 14).
1.1 Contributions
The following are the main contributions of the paper.
  • Construction of initial algebras for the non-polynomial functors $\operatorname{P}^{n}_{U}$, generalising the construction of the type of iterative of sets as the initial algebra for the powerset functor to higher type levels (Theorem 15).
  • Proof that these initial algebras are $n$-type universes of $n$-types, with definitional decoding (Section 7).
  • Generalisation of the axioms of set theory to properties of $\in$-structures of any type level (Section 2).
  • A framework for representations of types in $\in$-structures. This is applied to give a new formulation of the axiom of infinity, which does not fix a specific encoding of the natural numbers (Definition 5.7).
  • Equivalence of $\operatorname{P}^{n}_{U}$-coalgebras and $U$-like $\in$-structures, generalising the well-known connection between coalgebra and set theory (Theorem 3).
  • Proof that any fixed-point of $\operatorname{P}^{n}_{U}$ is a model of the generalisations of the axioms of set theory, except foundation, both generalising and proving in HoTT the result by Rieger [rieger1957] (Section 5).
  • New and short proof of the fiberwise equivalence lemmas: equivalence of families of maps (resp. equivalences) and maps (resp. equivalences) of total spaces respecting the first coordinate (Lemma 5 and Corollary 2).
Some of the ideas and definitions of this article were present in an unpublished preprint, titled “Non-wellfounded sets in HoTT” [GylterudBonnevier2020]. This preprint however, had a flawed argument in its fourth section and the main construction of that preprint cannot be carried out as described there. The results from Sections 2 and 3 of the preprint, which were correct, have been generalised to higher type levels. These generalisations can now be found in Section 2 and 5 of the current paper.
1.2 Formalisation
Everything in this paper has been formalised in the Agda proof assistant [agda]. Our formalisation builds on the agda-unimath library [agda-unimath], which is an extensive library of formalised mathematics from the univalent point of view.
The formalisation for this paper can be found at: https://git.app.uib.no/hott/hott-set-theory. Throughout the paper there will also be clickable links to specific lines of Agda code corresponding to a given result. These will be shown as the Agda logo .
1.3 Notation and universes
A lot of the basic constructions within HoTT have an established notation at this point in time. Nevertheless, to avoid confusion, we include here a list of some of the, perhaps less obvious, notation we will use in this paper. The notation we do not include in this list will usually follow the conventions of the HoTT Book [hottbook].
  • $\mathbf{0}$ denotes the empty type, with eliminator $\operatorname{ex\mbox{-}falso}$.
  • $\mathbf{1}$ denotes the unit type.
  • $\mathbf{2}$ denotes the type with two elements.
  • $\mathbb{N}$ denotes the type of natural numbers, with constructors 00 and $\operatorname{s}$.
  • $\operatorname{id\mbox{-}equiv}$ denotes the identity equivalence, on a given type.
  • $f\sim g$ denotes the type of homotopies from $f$ to $g$: $\prod_{x:X}f\,x=g\,x$.
  • $\operatorname{refl\mbox{-}htpy}$ denotes the homotopy $f\sim f$ given by the map $\lambda\,x.\,\operatorname{refl}$.
  • $\operatorname{fiber}\,f\,y$ denotes the homotopy fiber: $\sum_{x:X}f\,x=y$.
  • $\pi_{0}$ and $\pi_{1}$ denote the first, respectively second, projection out of a $\Sigma$-type.
  • Given a path $p:x=y$, $p^{-1}$ denotes the inverse path $y=x$.
  • Given a family $P$ of types over $X$ and a path $p:x=y$, $\operatorname{tr}^{P}_{p}$ denotes the transport function from $P\,x$ to $P\,y$ over $p$.
  • Given types $A$ and $B$ and a path $p:A=B$, $\operatorname{coe}\,p:A\rightarrow B$ is the map defined by path induction, taking the identity map for $\operatorname{refl}$.
  • Given an invertible function $f$ (usually an equivalence), $f^{-1}$ denotes the inverse.
  • Given a family of maps $f:\prod_{x:X}P\,x\rightarrow Q\,x$, $\operatorname{tot}\,f:\sum_{x:X}P\,x\rightarrow\sum_{x:X}Q\,x$ is the function: $\lambda\,(x,p).\,(x,f\,x\,p)$.
  • $\operatorname{is\mbox{-}}n\operatorname{\mbox{-}trunc\mbox{-}map\,}f$ is the proposition that $f$ is an $n$-truncated map: $\prod_{y:Y}\operatorname{is\mbox{-}}n\operatorname{\mbox{-}type\,}(\operatorname{fiber}\,f\,y)$.
  • $X\hookrightarrow Y$ is the type of propositionally truncated maps: $\sum_{f:X\rightarrow Y}\operatorname{is\mbox{-}}(-1)\operatorname{\mbox{-}trunc\mbox{-}map\,}f$.
  • $X\hookrightarrow_{n}Y$ is the type of $n$-truncated maps: $\sum_{f:X\rightarrow Y}\operatorname{is\mbox{-}}n\operatorname{\mbox{-}trunc\mbox{-}map\,}f$.
  • $X\twoheadrightarrow Y$ is the type of $(-1)$-connected maps: $\sum_{f:X\rightarrow Y}\prod_{y:Y}\operatorname{is\mbox{-}contr\,}\|\operatorname{fiber}\,f\,y\|_{-1}$.
  • $X\twoheadrightarrow_{n}Y$ is the type of $n$-connected maps: $\sum_{f:X\rightarrow Y}\prod_{y:Y}\operatorname{is\mbox{-}contr\,}\|\operatorname{fiber}\,f\,y\|_{n}$
  • $\exists!_{x:X}P\,x$ denotes the type: $\operatorname{is\mbox{-}contr\,}\left(\sum_{x:X}P\,x\right)$.
  • $\operatorname{funext}$ is the function $f\sim g\rightarrow f=g$ given by function extensionality.
  • $\operatorname{ua}$ is the function $X\simeq Y\rightarrow X=Y$ given by univalence.
  • $\operatorname{Prop}_{U}$ is the type of all propositions in the universe $U$, i.e. the type $\sum_{X:U}\operatorname{is\mbox{-}prop\,}X$.
  • $\operatorname{Set}_{U}$ is the type of all sets in the universe $U$, i.e. the type $\sum_{X:U}\operatorname{is\mbox{-}set\,}X$.
  • More generally, $n\operatorname{\mbox{-}Type}_{U}$ is the type of all $n$-types in the universe $U$, i.e. the type $\sum_{X:U}\operatorname{is\mbox{-}}n\operatorname{\mbox{-}type\,}X$.
We will use the same terminology as the HoTT Book regarding type levels. But we will also define a notion of level for $\in$-structures and elements in (the carrier of) an $\in$-structure. This overloading of terminology should be fine however, since it should be clear from the context what kind of level we are referring to. The notions of mere proposition and mere set are used to denote types of level $-1$ and 00 respectively, when there is need for clarity.
In this paper we will assume two type universes***The formalisation differs in this respect as it uses a hierarchy of universes, but the relationship between $U$ and $\operatorname{Type}$ in the article, is the same as the relationship between UU i and UU (lsuc i) in the formalisation., a large univalent universe, denoted $\operatorname{Type}$, and a small univalent universe, denoted $U$. We use cumulative universes, i.e. $U:\operatorname{Type}$ and $X:\operatorname{Type}$, for all $X:U$. It is assumed that both $U$ and $\operatorname{Type}$ are closed under the usual type formers, such as $\Pi$-types, $\Sigma$-types, and identity types. The constructions below also use the empty type and the type of natural numbers. We will use function extensionality freely.
From Section 4 and onwards, we will also assume that we can construct small images in certain situations. This assumption is informed by Rijke’s modified join construction [rijke2017], which can be used to construct such small images. One can alternatively assume that $U$ is closed under homotopy colimits, from which the smallness of (certain) images follows by the join construction.
2 $\in$-structures
In this section we give the definition of $\in$-structuresThis section defines the notion of $\in$-structure slightly differently than $\in$-structures were defined in a previous article by one of the authors [gylterud-iterative]. This difference is by incorporating extensionality and generalising to higher type levels. and formulate properties of these inspired by set theory. Most of the properties are indexed by a type level, from 0 to $\infty$. The level 0 version of the property is equivalent to the usual set theoretic concept for $\in$-structures of level 0, while the $\infty$ version was explored in [gylterud-multisets].
Definition 1 ( ).
An $\in$-structure is a pair $(V,\in)$ where $V:\operatorname{Type}$ and $\in\,:V\rightarrow V\rightarrow\operatorname{Type}$, which is extensional: for each $x,y:V$, the canonical map $x=y\rightarrow\prod_{z:V}{z\in x}\simeq{z\in y}$ is an equivalence of types.
Extensionality states that we can distinguish sets by their elements. It is expressed in first order logic using logical equivalence, but since we are working in the framework of HoTT and want to allow for elementhood relations of higher type level, we use instead equivalence of types. This of course reduces to logical equivalence in the case when the $\in$-relation is propositional.
Many times we will want to talk about all members of a given element in $V$. We introduce a notation for this.
Definition 2 ( ).
Given an $\in$-structure, $(V,\in)$, we define the family $\operatorname{El}:V\rightarrow\operatorname{Type}$ by $\operatorname{El}\,a:=\sum_{x:V}x\in a$.
The usual notion of (extensional) model of set theory corresponds to $\in$-structures, $(V,\in)$, where $x\in y$ is a mere proposition for each $x,y:V$ (and consequently $V$ is a mere set). However, there are examples of extensional $\in$-structures where $V$ is not a mere set. One such example, based on Aczel’s [aczel1978] type $\operatorname{W}_{A:U}A$, was explored in an article by one of the authors [gylterud-multisets].
We can stratify $\in$-structures based on the type level of the $\in$-relation.
Definition 3 ( ).
Given $n:\mathbb{N}_{-2}$, an $\in$-structure $(V,\in)$ is said to be of level (n+1) if for every $x,y:V$ the type $x\in y$ is an $n$-type.
The following proposition explains the occurrence of $n+1$ in the definition above:
Proposition 1 ( ).
In an $\in$-structure, $(V,\in)$, of level $n$ the type $V$ is an $n$-type.
Proof.
Let $(V,\in)$ be an $\in$-structure of level $(n+1)$, for $n:\mathbb{N}_{-2}$. By extensionality $x=y$ is equivalent to $\prod_{z:V}{z\in x}\simeq{z\in y}$, which is an $n$-type, hence $V$ is an $(n+1)$-type.∎
Remark: By definition there are no $-2$ level $\in$-structures, and the $-1$ level structures all have trivial $\in$-structure. Thus, we shall from here on focus on the $\in$-structures of level 0 or higher.
Of special interest will be the elements of an $\in$-structure which look like sets, in the sense that elementhood is a proposition. This is by definition the case for all elements in $\in$-structures of level 0, but such elements may occur in structures of all levels.
Definition 4 ( ).
An element $x:V$ is a $(k+1)$-type in $(V,\in)$ if $y\in x$ is of level $k:\mathbb{N}_{-1}$ for all $y:V$.
If $x:V$ is a 00-type in $(V,\in)$, we also say it is a mere set in $(V,\in)$.
Proposition 2 ( ).
If $x:V$ is a $(k+1)$-type in $(V,\in)$ then $x=y$ is a $k$-type for any $y:V$.
Proof.
By extensionality, the type $x=y$ has the same type level as the type $\prod_{z:V}z\in x\simeq z\in y$. The latter is a $k$-type as $z\in x$ is a $k$-type for every $z:V$.∎
Note that $x:V$ being a $k$-type in $(V,\in)$ does not imply that $\operatorname{El}\,x$ is a $k$-type. However, if $V$ is a $k$-type then $\operatorname{El}\,x$ is a $k$-type if $x$ is a $k$-type.
2.1 Ordered pairing
The characteristic property of ordered pairs is that two pairs are equal if and only if the first coordinates are equal and the second coordinates are equal, i.e. it is a pair where the order of the elements matters. In our setting this means that the ordered pair of two elements $x,y:V$ should be an element $\langle x,y\rangle:V$ such that for any other ordered pair $\langle x^{\prime},y^{\prime}\rangle$, for some $x^{\prime},y^{\prime}:V$, $\langle x,y\rangle=\langle x^{\prime},y^{\prime}\rangle$ exactly when $x=x^{\prime}$ and $y=y^{\prime}$. Using the characterisation of the identity types of cartesian products, this is equivalently saying that $\langle x,y\rangle=\langle x^{\prime},y^{\prime}\rangle$ exactly when $(x,y)=_{V\times V}(x^{\prime},y^{\prime})$. In order to allow for higher level $\in$-structures the “exactly when” should be replaced with type equivalence. Moreover, we want it to be the canonical one, in the sense that if $x\equiv x^{\prime}$ and $y\equiv y^{\prime}$ then the equivalence should send $\operatorname{refl}$ to $\operatorname{refl}$.
In HoTT, this is the statement that ordered pairing is an embedding $V\times V\hookrightarrow V$. This neatly encapsulates and generalises the usual characterisation of equality of ordered pairs in a way that is completely independent of the level of the $\in$-structure. But it does not uniquely define the encoding of ordered pairs. Indeed, there are several ways to encode ordered pairs in ordinary set theory. The usual Kuratowsky pairing will work for 0-level $\in$-structures, but not for higher level structures. Luckily, Norbert Wiener’s encoding, $\langle x,y\rangle:=\{\{\{x\},\emptyset\},\{\{y\}\}\}$, will work for structures of any level.
This was originally observed in previous work by one of the authors [gylterud-multisets]. However, there the equivalence $\left(\langle x,y\rangle=\langle x^{\prime},y^{\prime}\rangle\right)\simeq\left((x,y)=_{V\times V}(x^{\prime},y^{\prime})\right)$ was not required to be the canonical one, as we require here. Moreover, the encoding of ordered pairs used the untruncated variants of singletons and unordered pairs. But it may be the case that we can only construct truncated variants in a given $\in$-structure. Therefore, we will here construct ordered pairs, based on the Wiener encoding, but for any truncation level of singletons and unordered pairs.
Since the results which follow are independent of encoding, we will not commit to any specific way of forming ordered pairs, but simply assume ordered pairing as an extra structure.
Definition 5 ( ).
Given an $\in$-structure, $(V,\in)$, an ordered pairing structure on $(V,\in)$ is an embedding $V\times V\hookrightarrow V$.
While having ordered pairing is a structure, once the pairing structure is fixed, the notion of being an ordered pair is a proposition. As is clear from the proof below, this fact is just another formulation of the characterisation of equality of ordered pairs.
Proposition 3.
Being an ordered pair is a mere proposition: for a fixed $\in$-structure $(V,\in)$ with ordered pairing structure $\langle$-,$\rangle$ - , - ⟩, the type $\sum_{a,b:V}\langle a,b\rangle=x$ is a proposition, for all $x:V$.
Proof.
Observe that $\sum_{a,b:V}\langle a,b\rangle=x$ is the fiber of $\langle$-,$\rangle$ - , - ⟩ over $x$, which is a proposition since $\langle$-,$\rangle$ - , - ⟩ is an embedding.∎
2.2 Properties of $\in$-structures
In this section we explore how further set-theoretic notions, such as pairing, union, replacement, separation and exponentiation, can be expressed as mere propositions about $\in$-structures. As mentioned in the introduction, these notions can be generalised in different ways by using different levels of truncation. When characterising unions, for instance, it makes a difference whether the truncated existential quantifier or dependent pair types are used. With the truncated existential quantifier we only get one copy of each element in the union, but with the dependent pair type we may get more copies.
A given $\in$-structure can satisfy several versions, but a recurring theme is that $n$-level structures will only satisfy the $k$-truncated versions for $k\leq n$.
Convention: In the rest of the paper we will consider the type of truncation levels to be the type $\mathbb{N}^{\infty}_{-2}$, i.e. the usual truncation levels extended by an element $\infty$ for which $\|P\|_{\infty}$ is defined as $P$ and such that $\infty-1=\infty=\infty+1$.
The properties of $\in$-structures will be parameterised by truncation level. If the truncation level is omitted, in the notation or reference to a property, we mean the variant that is labeled with 00 (which usually involves $(-1)$-truncation in the definition).
Interestingly, even the untruncated versions of the set theoretic properties end up being propositions. For instance, to have all $\infty$-unions is a mere property of $\in$-structures. This is because the properties characterise the material sets they claim existence of up to equality, by extensionality.
Proposition 4 ( ).
Given an $\in$-structure, $(V,\in)$, let $\phi:V\rightarrow\operatorname{Type}$ be a type family on $V$. Then the type $\sum_{x:V}\prod_{z:V}z\in x\simeq\phi\,z$ is a proposition.
Proof.
Assume $(x,\alpha):\sum_{x:V}\prod_{z:V}z\in x\simeq\phi\,z$, then it is enough to show that the type is contractible. We have the following chain of equivalences:
$\left(\sum_{x^{\prime}:V}\prod_{z:V}z\in x^{\prime}\simeq\phi\,z\right)\simeq\left(\sum_{x^{\prime}:V}\prod_{z:V}z\in x^{\prime}\simeq z\in x\right)\simeq\left(\sum_{x^{\prime}:V}x^{\prime}=x\right)$
The last type is contractible.∎
Remark: Proposition 4 states that the generalisation of the unrestricted comprehension, $\{z|\phi\,z\}$, determines a set uniquely, when existent. Many existence statements in set theory can be seen as fleshing out for which forms $\phi$ this comprehension defines a set.
2.2.1 Unordered tuples
The usual notion of pairing naturally extends to unordered tupling of any arity. The arity can be expressed by any type. The usual pairing operation is tupling with respect to the booleans, the singleton operation is tupling for the unit type, and the empty set is tupling for the empty type. We generalise the property of having tupling to all truncation levels.
Definition 6 ( ).
Given $k:\mathbb{N}^{\infty}_{-1}$ and a type $I$, an $\in$-structure, $(V,\in)$, has $(k+1)$-unordered $I$-tupling if for every $v:I\rightarrow V$ there is $\{v\}_{k+1}:V$ such that $\prod_{z:V}z\in\{v\}_{k+1}\simeq\|\sum_{i:I}v\,i=z\|_{k}$
For the special cases $I=\operatorname{Fin}\,n$, we say that $(V,\in)$ has $k$-unordered $n$-tupling. If $(V,\in)$ has $k$-unordered $n$-tupling for every $n:\mathbb{N}$, we say that $(V,\in)$ has finite, $k$-unordered tupling.
We will use the usual notation for finite tuplings, but with a subscript for the truncation level: $k$-unordered $n$-tupling is denoted by $\{x_{0},\cdots,x_{n-1}\}_{k}$. For $k$-unordered $\mathbf{0}$-tupling we will use the notation $\emptyset$. Observe that $\{x_{0},\cdots,x_{n-1}\}_{k}$ is a $k$-type in $(V,\in)$.
The set $\emptyset$ is the set with no elements. For any $z:V$ we have:
$z\in\emptyset\simeq\Big{\|}\sum_{i:\mathbf{0}}\operatorname{ex\mbox{-}falso}\,i=z\ \Big{\|}_{k-1}\simeq\|\ \mathbf{0}\ \|_{k-1}\simeq\mathbf{0}$ (1)
Note here that the truncation level $k$ does not matter (hence why we exclude it from the notation $\emptyset$). By extensionality, the sets corresponding to $k$-unordered and $k^{\prime}$-unordered $\mathbf{0}$-tupling, for any two $k$ and $k^{\prime}$, are equal. Note also that $\emptyset$ is a mere set in any $\in$-structure.
Remark: We say that an $\in$-structure, $(V,\in)$, has empty set if there is an element $x:V$ such that $\prod_{z:V}z\in x\simeq\mathbf{0}$. By the previous paragraph, this is equivalent to saying that $(V,\in)$ has $\mathbf{0}$-tupling.
However, the singletons $\{x\}_{k}$ may be different for different truncation levels. For any $x,z:V$ we have:
$z\in\{x\}_{k}\simeq\Big{\|}\sum_{i:\mathbf{1}}x=z\ \Big{\|}_{k-1}\simeq\|\ x=z\ \|_{k-1}$ (2)
In the special case when $x$ is a mere set in $(V,\in)$, the truncation level does not matter in that $\{x\}_{k}=\{x\}_{0}$, by extensionality. As an example: {}$\{\emptyset\}_{k}=\{\emptyset\}_{k^{\prime}}$ for any two $k$ and $k^{\prime}$, since $\emptyset$ is a mere set. Starting at level 1, repetitions matter in tupling: Given $k>0$ we have {,}$\{\emptyset,\emptyset\}_{k}\neq\{\emptyset\}_{k}$, since {,}$\emptyset\in\{\emptyset,\emptyset\}_{k}\simeq\|{\emptyset=\emptyset}+{\emptyset=\emptyset}\|_{k}\simeq 2$ while {}$\emptyset\in\{\emptyset\}_{k}\simeq\|\emptyset=\emptyset\|_{k}\simeq 1$.
An interesting usage of 0-unordered 1-tupling combined with 1-unordered $n$-tupling is the construction of a set $s_{n}$ such that $\operatorname{El}\,s_{n}=B(S_{n})$, the classifying type of the symmetric group on $n$ elements. Simply let $s_{n}=\{\{n\emptyset\}_{1}\}_{0}$, where $n\emptyset:V^{n}$ is the vector with $n$ copies of $\emptyset$.
2.2.2 Ordered pairs from unordered tuples
[…]
If $(V,\in)$ has level $n$, then we know that $(x,y)=(x^{\prime},y^{\prime})$ has type level $n-1$. For that case the $(n-1)$-truncated variants would give us the correct type level for $\langle x,y\rangle=\langle x^{\prime},y^{\prime}\rangle$. But, we observe that {}$\{-\}_{n}=\{-\}_{\infty}$ in $\in$-structures of level $n$ since we have $z\in\{x\}_{n}\simeq\|z=x\|_{n-1}\simeq(z=x)$. For unordered pairs we need to distinguish between the $(-1)$-truncated case and all other cases since coproducts are not closed under propositions. More specifically, for $n\geq 1$ we have {,}$\{-,-\}_{n}=\{-,-\}_{\infty}$ if $(V,\in)$ has level $n$, since $z\in\{x,y\}_{n}\simeq\|(z=x)+(z=y)\|_{n-1}\simeq((z=x)+(z=y))$. This equivalence does not hold for $n=0$ and arbitrary $x,y:V$. However, if $x\neq y$, then $z=x\rightarrow z\neq y$ and the equivalence holds. We use these observations to make a general construction of ordered pairs which we can instantiate for $\in$-structures of all levels.
Lemma 1 ( ).
Given an $\in$-structure, $(V,\in)$, with an operation $\alpha:V\times V\rightarrow V$ with equivalences $e:\prod_{x,y:V}x\neq y\rightarrow\prod_{z:V}z\in\alpha(x,y)\simeq((z=x)+(z=y))$, and two disjoint embeddings, $f\ g:V\hookrightarrow V$, i.e. $\prod_{x,y:V}f\,x\neq g\,y$, then $\alpha\circ(f\times g):V\times V\rightarrow V$ is an embedding.
Proof.
We need to show, for any $(x,y),(x^{\prime},y^{\prime}):V\times V$, that $\operatorname{ap}_{\alpha\circ(f\times g)}:(x,y)=(x^{\prime},y^{\prime})\rightarrow\alpha(f\,x,g\,y)=\alpha(f\,x^{\prime},g\,y^{\prime})$ is an equivalence. To this end, it is enough to construct some equivalence between the identity types that sends $\operatorname{refl}:\alpha(f\,x,g\,y)=\alpha(f\,x,g\,y)$ to $\operatorname{refl}:(x,y)=(x,y)$.
We have the following chain of equivalences:
$\left(\alpha(f\,x,g\,y)=\alpha(f\,x^{\prime},g\,y^{\prime})\right)$ $\simeq\prod_{z:V}z\in\alpha(f\,x,g\,y)\simeq z\in\alpha(f\,x^{\prime},g\,y^{\prime})$ (3)
$\simeq\prod_{z:V}((z=f\,x)+(z=g\,y))\simeq((z=f\,x^{\prime})+(z=g\,y^{\prime}))$ (4)
$\simeq\prod_{z:V}((z=f\,x)\simeq(z=f\,x^{\prime}))\times((z=g\,y)\simeq(z=g\,y^{\prime}))$ (5)
$\simeq\left(\prod_{z:V}(z=f\,x)\simeq(z=f\,x^{\prime})\right)\times\left(\prod_{z:V}(z=g\,y)\simeq(z=g\,y^{\prime})\right)$ (6)
$\simeq\left(f\,x=f\,x^{\prime}\right)\times\left(g\,y=g\,y^{\prime}\right)$ (7)
$\simeq\left(x=x^{\prime}\right)\times\left(y=y^{\prime}\right)$ (8)
$\simeq((x,y)=(x^{\prime},y^{\prime}))$ (9)
In step (3) we use extensionality for $(V,\in)$. In step (4) we use the equivalences $e\,(f\,x)\,(g\,y)$ and $e\,(f\,x^{\prime})\,(g\,y^{\prime})$, together with the fact that $f\,x\neq g\,y$ and $f\,x^{\prime}\neq g\,y^{\prime}$. The equivalence (5) follows from the fact that $z=f\,x$ and $z=g\,y^{\prime}$, and $z=f\,x^{\prime}$ and $z=g\,y$, are, respectively, mutually exclusive. In step (8) we use the fact that $f$ and $g$ are embeddings.
We chase $\operatorname{refl}:\alpha(f\,x,g\,y)=\alpha(f\,x,g\,y)$ through the equivalence:
$\operatorname{refl}$ $\mapsto\lambda\,z.\operatorname{id\mbox{-}equiv}$ (10)
$\mapsto\lambda\,z.\left(e\,z\right)\circ\operatorname{id\mbox{-}equiv}\circ\left(e\,z\right)^{-1}$ (11)
$=\ \lambda\,z.\operatorname{id\mbox{-}equiv}$ (12)
$\mapsto\lambda\,z.(\operatorname{id\mbox{-}equiv},\operatorname{id\mbox{-}equiv})$ (13)
$\mapsto(\lambda\,z.\operatorname{id\mbox{-}equiv},\lambda\,z.\operatorname{id\mbox{-}equiv})$ (14)
$\mapsto(\operatorname{refl},\operatorname{refl})$ (15)
$=\ \left(\operatorname{ap}_{f}\,\operatorname{refl},\operatorname{ap}_{g}\,\operatorname{refl}\right)$ (16)
$\mapsto\left(\operatorname{ap}_{f}^{-1}\left(\operatorname{ap}_{f}\,\operatorname{refl}\right),\operatorname{ap}_{g}^{-1}\left(\operatorname{ap}_{g}\,\operatorname{refl}\right)\right)$ (17)
$=\ (\operatorname{refl},\operatorname{refl})$ (18)
$\mapsto\ \operatorname{refl}$ (19)
where we have used the fact that extensionality for $(V,\in)$ sends $\operatorname{refl}$ to $\operatorname{id\mbox{-}equiv}$.∎
For the Norbert Wiener construction of ordered pairs we thus have to show that both {{}$\{\{-\}_{n}\}_{n}$ and {{}$\{\{-\}_{n},\emptyset\}_{n}$ are embeddings. Using the previous observation about the relationship between the $n$-truncated versions and the $\infty$-truncated ones, it is enough to show this for the $\infty$-truncated versions.
Lemma 2 ( ).
The function {}$\{-\}_{\infty}:V\rightarrow V$ is an embedding.
Proof.
We follow the same strategy as in the proof of Lemma 1. For any $x,y:V$ we have the following chain of equivalences:
($\left(\{x\}_{\infty}=\{y\}_{\infty}\right)$ $\simeq\prod_{z:V}z\in\{x\}_{\infty}\simeq z\in\{y\}_{\infty}$ (20)
$\simeq\prod_{z:V}\left(z=x\right)\simeq\left(z=y\right)$ (21)
$\simeq\left(x=y\right)$ (22)
In step (20) we use extensionality for $(V,\in)$.
Let $e:\prod_{x:V}\prod_{z:V}z\in\{x\}_{\infty}\simeq\left(z=x\right)$ be the defining family of equivalences for {$\{-\}_{\infty}$. We chase $\operatorname{refl}:\{x\}_{\infty}=\{x\}_{\infty}$ along the chain of equivalences above:
$\operatorname{refl}$ $\mapsto\lambda\,z.\operatorname{id\mbox{-}equiv}$ (23)
$\mapsto\lambda\,z.(e\,x\,z)\circ\operatorname{id\mbox{-}equiv}\circ(e\,x\,z)^{-1}$ (24)
$=\ \lambda\,z.\operatorname{id\mbox{-}equiv}$ (25)
$\mapsto\operatorname{refl}$ (26)
In step (23) we use the fact that extensionality for $(V,\in)$ sends $\operatorname{refl}$ to $\operatorname{id\mbox{-}equiv}$.∎
Lemma 3 ( ).
Let $\alpha:V\rightarrow V\rightarrow V$ be such that $e:\prod_{x,y:V}x\neq y\rightarrow\prod_{z:V}z\in\alpha\,x\,y\simeq((z=x)+(z=y))$. Then $\lambda\,x.\,\alpha\,\{x\}_{\infty}\,\emptyset:V\rightarrow V$ is an embedding.
Proof.
First, we observe that for all $x:V$, $\{x\}_{\infty}\neq\emptyset$ since $x\in\{x\}_{\infty}$ is inhabited but $x\in\emptyset$ is empty. We now follow the same strategy as in the proof of Lemma 1. For any $x,y:V$ we have the following chain of equivalences:
$\left(\alpha\,\{x\}_{\infty}\,\emptyset=\alpha\,\{y\}_{\infty}\,\emptyset\right)$ $\simeq\prod_{z:V}(z\in\alpha\,\{x\}_{\infty}\,\emptyset)\simeq(z\in\alpha\,\{y\}_{\infty}\,\emptyset)$ (27)
$\simeq\prod_{z:V}\left((z=\{x\}_{\infty})+(z=\emptyset)\right)\simeq\left((z=\{y\}_{\infty})+(z=\emptyset)\right)$ (28)
$\simeq\prod_{z:V}\left((z=\{x\}_{\infty})\simeq(z=\{y\}_{\infty})\right)\times\left((z=\emptyset)\simeq(z=\emptyset)\right)$ (29)
$\simeq\prod_{z:V}(z=\{x\}_{\infty})\simeq(z=\{y\}_{\infty})$ (30)
($\simeq\left(\{x\}_{\infty}=\{y\}_{\infty}\right)$ (31)
$\simeq\left(x=y\right)$ (32)
In step (27) we use extensionality for $(V,\in)$. In step (29) we use the fact that $\{x\}_{\infty}\neq\emptyset$ and $\{y\}_{\infty}\neq\emptyset$. In step (30) we use the fact that $(z=\emptyset)\simeq(z=\emptyset)$ is contractible since $\emptyset$ is a mere set, and hence $z=\emptyset$ is a proposition. In step (32) we use Lemma 2.
We chase $\operatorname{refl}:\alpha\,\{x\}_{\infty}\,\emptyset=\alpha\,\{x\}_{\infty}\,\emptyset$ along the chain of equivalences above:
$\operatorname{refl}$ $\mapsto\lambda\,z.\operatorname{id\mbox{-}equiv}$ (33)
$\mapsto\lambda\,z.(e\,z)\circ\operatorname{id\mbox{-}equiv}\circ(e\,z)^{-1}$ (34)
$=\ \lambda\,z.\operatorname{id\mbox{-}equiv}$ (35)
$\mapsto\lambda\,z.(\operatorname{id\mbox{-}equiv},\operatorname{id\mbox{-}equiv})$ (36)
$\mapsto\lambda\,z.\operatorname{id\mbox{-}equiv}$ (37)
$\mapsto\operatorname{refl}$ (38)
$=\ \operatorname{ap}_{\{-\}_{\infty}}\,\operatorname{refl}$ (39)
$\mapsto\operatorname{ap}_{\{-\}_{\infty}}^{-1}\,\left(\operatorname{ap}_{\{-\}_{\infty}}\,\operatorname{refl}\right)$ (40)
$=\ \operatorname{refl}$ (41)
In step (33) we use the fact that extensionality for $(V,\in)$ sends $\operatorname{refl}$ to $\operatorname{id\mbox{-}equiv}$.∎
Theorem 1 ( ).
[…]
Proof.
Corollary of Lemma 1, Lemma 2 and Lemma 3.∎
2.2.3 Restricted separation
In constructive set theory [aczel1978], restricted separation is the ability to construct sets of the form $\{z\in x\,|\,\Phi\,z\}$, for a formula $\Phi$ where all quantifiers are bounded (i.e. $\forall a\in b\cdots$ and $\exists a\in b\cdots$). One way to internalise this to type theory is to require that the predicate $\Phi:V\rightarrow\operatorname{Prop}_{U}$ is a predicate of propositions in $U$. We do the same here in that we require the predicate to take values in $U$, but as with the other properties, we generalise to higher truncation levels.
Definition 7 ( ).
[…]
Remark: Usually, in set theory, the predicate is defined for any $x$, even when taking the restricted separation. But this falls under the above: Assume $Q:V\rightarrow k\operatorname{\mbox{-}Type}_{U}$, then one can readily define $P(z,e)=Q\,z$ and the defining property becomes $z\in\{x\,|\,P\}\simeq(z\in x\times Q\,z)$. However, going in the opposite direction only works at level 0, because the predicate $P$ may otherwise depend on the specific witness of elementhood.
Remark: At level 1, we can use 00-unordered $1$-tupling, 1-unordered tupling and $U$-restricted $1$-separation, to construct for every group $G$ a set $s_{G}$ such that $\operatorname{El}\,s_{G}=B(G)$, the classifying type for the group.For a thorough treatment of groups and classifying types, see the Symmetry book [bezem2022symmetry] First let $|G|\emptyset:|G|\rightarrow V$ be the constant function mapping every element of the group to $\emptyset$. This is the tuple to which we will apply 1-unordered tupling, constructing $\{|G|\emptyset\}_{1}$. This multiset has one copy of $\emptyset$ for every element of the group, hence its automorphisms in $V$ are the bijections on $|G|$:
($\left(\{|G|\emptyset\}_{1}=\{|G|\emptyset\}_{1}\right)$ $\simeq\prod_{z:V}(z\in\{|G|\emptyset\}_{1})\simeq(z\in\{|G|\emptyset\}_{1})$ (42)
$\simeq\prod_{z:V}(|G|\times(z=\emptyset))\simeq(|G|\times(z=\emptyset))$ (43)
$\simeq\left(|G|=_{U}|G|\right)$ (44)
If we take the 0-singleton $\{\{|G|\emptyset\}_{1}\}_{0}$, we get a set which has $B(\operatorname{Aut}_{U}\,|G|)$ as its type of elements:
$\operatorname{El}\,\{\{|G|\emptyset\}_{1}\}_{0}$ $=\sum_{z:V}z\in\{\{|G|\emptyset\}_{1}\}_{0}$ (45)
$\simeq\sum_{z:V}\|z=\{|G|\emptyset\}_{1}\|_{-1}$ (46)
$=B(\operatorname{Aut}_{V}\,\{|G|\emptyset\}_{1})$ (47)
$\simeq B(\operatorname{Aut}_{U}\,|G|)$ (48)
The final ingredient is that we need a map $f:B(G)\rightarrow B(\operatorname{Aut}\,|G|)$. This map is the well-known map induced by multiplication in the group. This is a cover, i.e. its fibers are sets, hence $\operatorname{fiber}\,f:B(\operatorname{Aut}\,|G|)\rightarrow 0\operatorname{\mbox{-}Type}$, which we will coerce along the equivalence $B(\operatorname{Aut}_{U}\,|G|)\simeq\operatorname{El}\,\{\{|G|\emptyset\}_{1}\}_{0}$ to obtain $P_{G}:\operatorname{El}\,\{\{|G|\emptyset\}_{1}\}_{0}\rightarrow 0\operatorname{\mbox{-}Type}$. Thus, we define $s_{G}:=\{\{\{|G|\emptyset\}_{1}\}_{0}|P_{G}\}$, which then has the desired property:
$\operatorname{El}\,s_{G}$ $=\sum_{z:V}z\in\{\{\{|G|\emptyset\}_{1}\}_{0}|P_{G}\}$ (49)
$\simeq\sum_{z:V}\sum_{e:z\in\{\{|G|\emptyset\}_{1}\}_{0}}P_{G}(z,e)$ (50)
$\simeq\sum_{q:\operatorname{El}\,\{\{|G|\emptyset\}_{1}\}_{0}}P_{G}\,q$ (51)
$\simeq\sum_{q:B(\operatorname{Aut}\,|G|)}\operatorname{fiber}\,f\,q$ (52)
$\simeq B(G)$ (53)
2.2.4 Replacement
Replacement says that given a set $a$, and a description of how to replace its elements, we can create a new set containing exactly the replacements of elements in $a$. Of course, two elements in $a$ may be replaced by the same elements, so the property must include truncation.
Definition 8 ( ).
An $\in$-structure, $(V,{\in})$, has $(k+1)$-replacement, for $k:\mathbb{N}^{\infty}_{-1}$, if for every $a:V$ and $f:\operatorname{El}\,a\rightarrow V$ there is an element $\{f(x)\,|\,x\in a\}:V$ such that $\prod_{z:V}z\in\{f(x)\,|\,x\in a\}\simeq\big{\|}\sum_{x:\operatorname{El}\,a}f\,x=z\;\big{\|}_{k}$
Remark: Having $k$-replacement is to have $k$-unordered $I$-tupling for $I=\operatorname{El}a$, for any $a:V$.
In Section 3 we further explore the implications of $k$-replacement and how it affects representations of types within $\in$-structures.
2.2.5 Union
Just as for tupling, there is a notion of union to consider for each type level, for $\in$-structures, distinguished by the truncation of the existential quantifier.
Definition 9 ( ).
An $\in$-structure, $(V,\in)$, has $(k+1)$-union, for $k:\mathbb{N}^{\infty}_{-1}$, if for every $x:V$ there is $\bigcup_{k}x:V$ such that $\prod_{z:V}z\in\bigcup_{k}x\simeq\big{\|}\sum_{y:V}\;z\in y\ \times\ y\in x\;\big{\|}_{k}$.
As usual, from pairing and general unions, one can define binary $k$-union by $x\cup_{k}y:=\bigcup_{k}\{x,y\}_{k}$. For $z:V$ we have:
$z\in x\cup_{k}y$ $\simeq\Big{\|}\sum_{w:V}\;z\in w\ \times\ w\in\{x,y\}_{k}\Big{\|}_{k-1}$ (54)
$\simeq\Big{\|}\sum_{w:V}\;z\in w\ \times\ \|w=x+w=y\|_{k-1}\Big{\|}_{k-1}$ (55)
$\simeq\Big{\|}\sum_{w:V}\;z\in w\ \times\ \left(w=x+w=y\right)\Big{\|}_{k-1}$ (56)
$\simeq\Big{\|}z\in x+z\in y\Big{\|}_{k-1}$ (57)
Binary 00-union is the usual binary union in set theory, where copies are discarded. The higher level binary union keeps copies. For example, {}$\{\emptyset\}_{1}\cup_{0}\{\emptyset\}_{1}=\{\emptyset\}_{0}$, while {}$\{\emptyset\}_{1}\cup_{k}\{\emptyset\}_{1}=\{\emptyset,\emptyset\}_{k}$, for $k\geq 1$.
2.2.6 Exponentiation
Exponentiation states that for any two sets $a$ and $b$ there is a set that contains exactly the functions from $a$ to $b$. In order to express this property we need to first generalise the notion of a function internal to an $\in$-structure. As opposed to the other properties, the notion of a function can be expressed in a uniform way for $\in$-structures of any level. As in usual set theory, the notion of a function is relative to a choice of ordered pairing structure.
But what constitutes a function between two generalised sets, say $a,b:V$? The perhaps easiest answer, which is the one we will argue for here, is that it is in essence a function $\operatorname{El}\,a\rightarrow\operatorname{El}\,b$, which can then be represented as set itself by taking its graph. At level 0, in usual set theory, a function is completely determined by its graph. The general situation is, however, that if $a$ and $b$ are of level $n+1$ there is an $n$-type of function structures which can be put on a set of pairs. We call this structure $\operatorname{operation}_{a\,b}$, and Proposition 8 constructs an equivalence between $\operatorname{El}\,a\rightarrow\operatorname{El}\,b$ and sets with $\operatorname{operation}_{a\,b}$-structure.
Remark: As usual in constructive set theory, we use the exponentiation axiom instead of the powerset axiom. One could define the powerset axiom for higher level structures by saying that $z:V$ is a subset of $x:V$ if for all $y:V$ there is an embedding $y\in z\hookrightarrow y\in x$. However, in any fixed-point model, the powerset axiom would require a small subobject classifier in the ambient type theory, hence why we do not consider it further in this paper.
Definition 10 ( ).
Given three elements $a,b,f:V$ in an $\in$-structure $(V,\in)$ with ordered pairing $\langle-,-\rangle:V\times V\hookrightarrow V$, define a type:
$\operatorname{operation}_{a\,b}\,f:=$ $\left(\prod_{x:V}\,x\in a\simeq\sum_{y:V}\langle x,y\rangle\in f\right)$
$\times$ $\left(\prod_{x:V}\prod_{y:V}\langle x,y\rangle\in f\rightarrow y\in b\right)$
$\times$ $\left(\prod_{z:V}\,z\in f\rightarrow\sum_{x:V}\,\sum_{y:V}\,z=\langle x,y\rangle\right)$
The first conjunct of $\operatorname{operation}_{a\,b}\,f$ states that the pairs in $f$ form an operation with domain $a$, and the second states that its codomain is $b$. The third conjunct, which is always a proposition by Proposition 3, states that $f$ only contains pairs.
Note that $\operatorname{operation}$ here is the same as Definition 8 in [gylterud-multisets]. However, in this paper we define another type which we show is equivalent to $\operatorname{operation}$ and which is of a more type theoretic flavor.
The type $\operatorname{operation}_{a\,b}\,f$ is not always a proposition. Hence, an element $p:\operatorname{operation}_{a\,b}\,f$ should be regarded as an operation structure on $f$ with domain $a$ and codomain $b$. However, in the case when $a$ and $b$ are mere sets, $\operatorname{operation}_{a\,b}\,f$ is a proposition.
Proposition 5 ( ).
Given three elements $a,b,f:V$ in an $\in$-structure $(V,\in)$ with ordered pairing $\langle-,-\rangle:V\times V\hookrightarrow V$, if $a$ and $b$ are mere sets in $(V,\in)$, then $\operatorname{operation}_{a\,b}\,f$ is a proposition equivalent to the following type:
$\left(\prod_{x:V}\,x\in a\rightarrow\underset{y:V}{\exists!}\langle x,y\rangle\in f\right)$
$\times$ $\left(\prod_{x:V}\prod_{y:V}\langle x,y\rangle\in f\ \rightarrow\ x\in a\ \times\ y\in b\right)$
$\times$ $\left(\prod_{z:V}\,z\in f\rightarrow\sum_{x:V}\,\sum_{y:V}\,z=\langle x,y\rangle\right)$
Proof.
We start by proving the equivalence. First, we observe that for any $x:V$, $p:x\in a$, and function $e:\sum_{y:V}\langle x,y\rangle\in f\rightarrow x\in a$ we have an equivalence
$\operatorname{fiber}\,e\ p\ \simeq\sum_{y:V}\langle x,y\rangle\in f$
since for any $q:\sum_{y:V}\langle x,y\rangle\in f$ the type $e\,q=p$ is contractible because $a$ is a mere set. It thus follows that for any $x:V$ we have
$\left(x\in a\simeq\sum_{y:V}\langle x,y\rangle\in f\right)$ $\simeq\sum_{e:\sum_{y:V}\langle x,y\rangle\in f\rightarrow x\in a}\prod_{p:x\in a}\operatorname{is\mbox{-}contr\,}(\operatorname{fiber}\,e\,p)$
$\simeq\left(\sum_{y:V}\langle x,y\rangle\in f\rightarrow x\in a\right)\times\left(x\in a\rightarrow\underset{y:V}{\exists!}\langle x,y\rangle\in f\right)$
The desired equivalence follows from this, after some currying and rearranging.
Since $\exists!_{y:V}\langle x,y\rangle\in f$ is a proposition it follows that $\left(\prod_{x:V}\,x\in a\rightarrow\exists!_{y:V}\langle x,y\rangle\in f\right)$ is a proposition. When $a$ and $b$ are mere sets, $x\in a\ \times\ y\in b$ is a proposition, and hence $\ \left(\prod_{x:V}\prod_{y:V}\langle x,y\rangle\in f\ \rightarrow\ (x\in a)\ \times\ (y\in b)\right)$ is a proposition. Finally, since ordered pairing is an embedding, and $\sum_{x:V}\,\sum_{y:V}\,z=\langle x,y\rangle$ is essentially the fibres of the pairing operation, it follows that
$\prod_{z:V}\,z\in f\rightarrow\sum_{x:V}\,\sum_{y:V}\,z=\langle x,y\rangle$
is a proposition.∎
This equivalence was proven as Lemma 6.10 in [gylterud-iterative] for the type $V$ constructed there. Here we show the equivalence for general $\in$-structures.
Corollary 1.
If $(V,\in)$ is of level 0 then $\operatorname{operation}_{a\,b}\,f$ is a proposition, for all $a,b,f:V$.
We can now define exponentiation for $\in$-structures.
Definition 11 ( ).
An $\in$-structure, $(V,\in)$ with an ordered pairing structure, has exponentiation if for every two $a,b:V$ there is an element $b^{a}$ such that $\prod_{f:V}(f\in b^{a})\simeq\operatorname{operation}_{a\,b}\,f$.
While $\operatorname{operation}$ is a straightforward internalisation of the set theoretic definition of a function, it can be inconvenient to work with. What follows is a definition that is equivalent to $\operatorname{operation}$ but which is sometimes easier to use.
Definition 12 ( ).
Given three elements $a,b,f:V$ in an $\in$-structure $(V,\in)$ with ordered pairing $\langle-,-\rangle:V\times V\hookrightarrow V$, define a type:
$\operatorname{operation^{\prime}}_{a\,b}\,f:=$ $\sum_{\phi:\operatorname{El}\,a\,\rightarrow\,\operatorname{El}\,b}\ \prod_{z:V}\left(z\in f\simeq\sum_{x:\operatorname{El}\,a}\langle\pi_{0}\ x,\pi_{0}\,(\phi\ x)\rangle=z\right)$
Proposition 6 ( ).
For any $a,b,f:V$ we have the following equivalence:
$\operatorname{operation}_{a\,b}\,f\simeq\operatorname{operation^{\prime}}_{a\,b}\,f$
Before we prove this equivalence we prove two general lemmas that we will need, concerning fiberwise functions and equivalences.
Lemma 4 ( ).
Let $A$ and $B$ be types and let $C:B\rightarrow\operatorname{Type}$ be a type family over $B$. For any embedding $f:A\hookrightarrow B$ and element $\gamma:\prod_{(b,c):\sum_{b:B}C\,b}\operatorname{fiber}\,f\ b$ we have an equivalence
$\sum_{a:A}C\,(f\ a)\simeq\sum_{b:B}C\,b.$
Proof.
Let $F:\sum_{a:A}C\,(f\ a)\rightarrow\sum_{b:B}C\,b$ be the function given by $F\,(a,c):=(f\,a,c)$. For $(b,c):\sum_{b:B}C\,b$, we have the following chain of equivalences:
$\operatorname{fiber}\,F\,(b,c)$ $\simeq\sum_{a:A}\sum_{c^{\prime}:C\,(f\,a)}\sum_{p:f\,a=b}\operatorname{tr}^{C}_{p}c^{\prime}=c$ (58)
$\simeq\sum_{a:A}\sum_{p:f\,a=b}\sum_{c^{\prime}:C\,(f\,a)}c^{\prime}=\operatorname{tr}^{C}_{p^{-1}}c$ (59)
$\simeq\operatorname{fiber}\,f\,b$ (60)
Step (60) uses the fact that $\sum_{c^{\prime}:C\,(f\,a)}c^{\prime}=\operatorname{tr}^{C}_{p^{-1}}c$ is contractible.
Thus, $\operatorname{fiber}\,F\ (b,c)$ is a proposition, since $f$ is an embedding, which is inhabited by $\gamma$, and therefore contractible.∎
There is an equivalence between fiberwise equivalences and equivalences of the total spaces that respect the first coordinate. This will be useful several times because it gives two equivalent characterisations of equality in slices over a type.
Lemma 5 ( ).
For any type $A$, and any families $P,Q:A\rightarrow\operatorname{Type}$ we have that
$\left(\prod_{x:A}P\,x\rightarrow Q\,x\right)\,\simeq\,\left(\sum_{\alpha:\sum_{x:A}P\,x\rightarrow\sum_{x:A}Q\,x}\pi_{0}\circ\alpha=\pi_{0}\right)$
Proof.
We have the following chain of equivalences:
$\left(\prod_{x:A}P\,x\rightarrow Q\,x\right)$ $\simeq\prod_{(x,\_):\sum_{x:A}P\,x}Q\,x$ (61)
$\simeq\prod_{(x,\_):\sum_{x:A}P\,x}\;\sum_{(x^{\prime},\_):\sum_{x^{\prime}:A}x^{\prime}=x}Q\,x^{\prime}$ (62)
$\simeq\prod_{(x,\_):\sum_{x:A}P\,x}\;\sum_{(x^{\prime},\_):\sum_{x^{\prime}:A}Q\,x^{\prime}}x^{\prime}=x$ (63)
$\simeq\sum_{\alpha:\sum_{x:A}P\,x\rightarrow\sum_{x:A}Q\,x}\pi_{0}\circ\alpha=\pi_{0}$ (64)
where (61) is currying, (62) follows from the fact that $\sum_{x^{\prime}:A}x^{\prime}=x$ is contractible and (64) is the interchange law between $\Sigma$-types and $\Pi$-types, together with function extensionality. The equivalence sends $f$ to $(\operatorname{tot}\,f,\operatorname{refl})$.∎
Remark: Lemma 5 has been independently added by Egbert Rijke to the agda-unimath library [agda-unimath]. The proof there is slightly different, with a direct construction of the equivalence, instead of equivalence reasoning.
Corollary 2 ( ).
For any type $A$, and any families $P,Q:A\rightarrow\operatorname{Type}$ we have an equivalence
$\left(\prod_{x:A}P\,x\,\simeq\,Q\,x\right)\,\simeq\,\left(\sum_{\alpha:\sum_{x:A}P\,x\,\simeq\,\sum_{x:A}Q\,x}\pi_{0}\circ\alpha=\pi_{0}\right)$
Proof.
This follows from Lemma 5 by an application of Theorem 4.7.7 in the HoTT Book [hottbook, p. 185], which states that a fiberwise transformation is a fiberwise equivalence if and only if the corresponding total function is an equivalence.∎
Proof of Proposition 6.
We outline the key steps of the equivalence. For full details see the Agda formalisation. We have the following chain of equivalences:
$\operatorname{operation}_{a\,b}\,f\simeq$ $\left(\prod_{z:\operatorname{El}\,f}\operatorname{fiber}\ \langle-,-\rangle\ (\pi_{0}\ z)\right)$ (65)
$\times$ $\left(\sum_{\sigma:\operatorname{El}\,a\simeq\sum_{(x,y):V\times V}\langle x,y\rangle\in f}\pi_{0}\sim\pi_{0}\circ\pi_{0}\circ\sigma\right)$
$\times$ $\left(\sum_{\phi:\sum_{(x,y):V\times V}\langle x,y\rangle\in f\rightarrow\operatorname{El}\,b}\pi_{0}\circ\phi\sim\pi_{1}\circ\pi_{0}\right)$
$\simeq$ $\sum_{\gamma:\prod_{z:\operatorname{El}\,f}\operatorname{fiber}\ \langle-,-\rangle\ (\pi_{0}\ z)}$ (66)
$\hskip 16.0pt\left(\sum_{\sigma:\operatorname{El}\,a\simeq\operatorname{El}\,f}\pi_{0}\circ\sigma^{-1}\sim\pi_{0}\circ\pi_{0}\circ\gamma\right)$
×$\hskip 6.0pt\times\left(\sum_{\phi:\operatorname{El}\,f\rightarrow\operatorname{El}\,b}\pi_{0}\circ\phi\sim\pi_{1}\circ\pi_{0}\circ\gamma\right)$
$\simeq$ $\sum_{\sigma:\operatorname{El}\,a\simeq\operatorname{El}\,f}\sum_{\phi:\operatorname{El}\,f\rightarrow\operatorname{El}\,b}\prod_{z:\operatorname{El}\,f}$ (67)
$\hskip 16.0pt\sum_{(x,y):V\times V}\left(\langle x,y\rangle=\pi_{0}\,z\right)\times\left(\pi_{0}\,(\sigma^{-1}\ z)=x\right)\times\left(\pi_{0}\,(\phi\ z)=y\right)$
$\simeq$ $\sum_{\sigma:\operatorname{El}\,a\simeq\operatorname{El}\,f}\sum_{\phi:\operatorname{El}\,f\rightarrow\operatorname{El}\,b}\prod_{z:\operatorname{El}\,f}\langle\pi_{0}\,(\sigma^{-1}\ z),\pi_{0}\,(\phi\ z)\rangle=\pi_{0}\ z$ (68)
$\simeq$ $\sum_{\phi:\operatorname{El}\,a\rightarrow\operatorname{El}\,b}\sum_{\sigma:\operatorname{El}\,a\simeq\operatorname{El}\,f}\prod_{x:\operatorname{El}\,a}\langle\pi_{0}\ x,\pi_{0}\,(\phi\ x)\rangle=\pi_{0}\,(\sigma\ x)$ (69)
$\simeq$ $\ \operatorname{operation^{\prime}}_{a\,b}\,f$ (70)
where in (65) we apply Corollary 2 to the first conjunct and Lemma 5 to the second conjunct, and then rearrange. In (66) we use Lemma 4 to construct an equivalence $\sum_{(x,y):V\times V}\langle x,y\rangle\in f\simeq\operatorname{El}\,f$ which we apply in the second and third conjuncts. In (67) we rearrange and then use the interchange law for $\prod$-types and $\sum$-types. In (68) we use extensionality for cartesian products and the fact that $\sum_{(x,y):V\times V}(\pi_{0}\,(\sigma^{-1}\ z),\pi_{0}\,(\phi\,z))=(x,y)$ is contractible. In (69) we swap $\operatorname{El}\,a$ for $\operatorname{El}\,f$ and then rearrange. Finally, in (70) we use Corollary 2 again.∎
The notion of $\operatorname{operation^{\prime}}$ captures type theoretic functions in the sense that the type of all operations from $a:V$ to $b:V$ is a subtype of the type $\operatorname{El}a\rightarrow\operatorname{El}b$.
Proposition 7 ( ).
Given $a,b:V$ in an $\in$-structure $(V,\in)$ with ordered pairing $\langle$-,$\rangle$ - , - ⟩, there is a canonical embedding
$\left(\sum_{f:V}\operatorname{operation^{\prime}}_{a\,b}\,f\right)\hookrightarrow\left(\operatorname{El}a\rightarrow\operatorname{El}b\right)$
Proof.
By swapping the $\Sigma$-types we have the following equivalence
$\left(\sum_{f:V}\operatorname{operation^{\prime}}_{a\,b}\,f\right)\simeq\sum_{\phi:\operatorname{El}a\rightarrow\operatorname{El}b}\sum_{f:V}\prod_{z:V}\left(z\in f\simeq\sum_{x:\operatorname{El}\,a}\langle\pi_{0}\ x,\pi_{0}\,(\phi\ x)\rangle=z\right)$ (71)
The type $\sum_{f:V}\prod_{z:V}\left(z\in f\simeq\sum_{x:\operatorname{El}\,a}\langle\pi_{0}\ x,\pi_{0}\,(\phi\ x)\rangle=z\right)$ is a proposition, by Proposition 4. The embedding is thus the composition of (71) with the first projection.∎
An operation from $a:V$ to $b:V$ is thus a function from $\operatorname{El}a\rightarrow\operatorname{El}b$. A natural question to ask is when all functions $\operatorname{El}a\rightarrow\operatorname{El}b$ correspond to some operation.
Proposition 8 ( ).
Given $a,b:V$ in an $\in$-structure $(V,\in)$ of level $n:\mathbb{N}^{\infty}$, with ordered pairing $\langle$-,$\rangle$ - , - ⟩, the embedding in Proposition 7 is an equivalence if $(V,\in)$ has $n$-replacement.
Proof.
Suppose $(V,\in)$ has $n$-replacement. By applying it to the map $\lambda\,x.\langle\pi_{0}\ x,\pi_{0}\,(\phi\ x)\rangle:\operatorname{El}a\rightarrow V$, we get an element of the following type:
$\sum_{f:V}\prod_{z:V}\left(z\in f\simeq\Big{\|}\sum_{x:\operatorname{El}\,a}\langle\pi_{0}\ x,\pi_{0}\,(\phi\ x)\rangle=z\Big{\|}_{n-1}\right)$
Claim: The type $\sum_{x:\operatorname{El}\,a}\langle\pi_{0}\ x,\pi_{0}\,(\phi\ x)\rangle=z$ is $(n-1)$-truncated.
Given that the claim is true we can drop the truncation. Thus the type
$\sum_{f:V}\prod_{z:V}\left(z\in f\simeq\sum_{x:\operatorname{El}\,a}\langle\pi_{0}\ x,\pi_{0}\,(\phi\ x)\rangle=z\right)$
is contractible, because it is an inhabited proposition, and hence the embedding constructed in Proposition 7 is an equivalence.
It remains to prove the claim. First, we have the following equivalence:
$\sum_{x:\operatorname{El}\,a}\left(\langle\pi_{0}\ x,\pi_{0}\,(\phi\ x)\rangle=z\right)\simeq\sum_{(x,y):\operatorname{El}a\times\operatorname{El}b}\left(\langle\pi_{0}\ x,\pi_{0}\,y\rangle=z\right)\times\left(\phi\ x=y\right)$ (72)
This follows from the fact that $\sum_{y:\operatorname{El}b}\phi\ x=y$ is contractible. The type $\phi\ x=y$ is $(n-1)$-truncated since $V$ is an $n$-type, by Proposition 1. The type $\sum_{(x,y):\operatorname{El}a\times\operatorname{El}b}\left(\langle\pi_{0}\ x,\pi_{0}\ y\rangle=z\right)$ is the fiber of the composite map $\langle-,-\rangle\circ(\pi_{0}\times\pi_{0})$ over $z:V$. The map $\langle$-,$\rangle$ - , - ⟩ is an embedding and thus an $(n-1)$-truncated map. The fibers of the two projection maps over $z$ are $z\in a$ and $z\in b$ respectively. These types are $(n-1)$-truncated. So the composite map is $(n-1)$-truncated, and the claim follows.∎
2.2.7 Accessible elements and foundation
We choose the same approach as the HoTT Book [hottbook] to well-foundedness, namely accessibility predicates. The axiom of foundation is then the statement that all elements are accessible. This need not be true in a given $\in$-structure, but the subtype of accessible elements inherits the $\in$-structure from the base type. This new $\in$-structure satisfies foundation.
Definition 13 ( ).
Given an $\in$-structure, $(V,\in)$, define inductively the predicate $\operatorname{Acc}:V\rightarrow\operatorname{Type}$ by
  • $\operatorname{acc}:\prod_{x:V}(\prod_{y:V}y\in x\rightarrow\operatorname{Acc}y)\rightarrow\operatorname{Acc}x$
Lemma 6 ( ).
For every $x:V$ the type $\operatorname{Acc}x$ is a mere proposition.
Proof.
Lemma 10.3.2 in the HoTT Book [hottbook, p. 454].∎
Definition 14 ( ).
An $\in$-structure, $(V,\in)$, has foundation if $\prod_{x:V}\operatorname{Acc}x$.
Since accessibility is a proposition, we can define for a given $\in$-structure, the subtype of accessible elements.
Definition 15 ( ).
Given an $\in$-structure, $(V,\in)$, define the type $V_{\operatorname{Acc}}:=\sum_{x:V}\operatorname{Acc}x$ and define the binary relation $\in_{\operatorname{Acc}}:=\lambda\,(x:V_{\operatorname{Acc}})\,(y:V_{\operatorname{Acc}}).\,\pi_{0}\,x\in\pi_{0}\,y$.
The subtype of accessible elements inherits the $\in$-structure from the base type.
Proposition 9 ( ).
Given an $\in$-structure, $(V,\in)$, the pair $(V_{\operatorname{Acc}},\in_{\operatorname{Acc}})$ forms an $\in$-structure.
Proof.
We need to prove extensionality for $(V_{\operatorname{Acc}},\in_{\operatorname{Acc}})$. For $x,y:V_{\operatorname{Acc}}$ we have the following chain of equivalences:
$\left(x=y\right)$ $\simeq\left(\pi_{0}\,x=\pi_{0}\,y\right)$ (73)
$\simeq\prod_{z:V}z\in\pi_{0}\,x\simeq z\in\pi_{0}\,y$ (74)
$\simeq\sum_{e\,:\,\sum_{z:V}z\in\pi_{0}\,x\ \simeq\ \sum_{z:V}z\in\pi_{0}\,y}\pi_{0}\circ e\sim\pi_{0}$ (75)
$\simeq\sum_{e\,:\,\sum_{z:V}\left(z\in\pi_{0}\,x\right)\times\left(\operatorname{Acc}z\right)\ \simeq\ \sum_{z:V}\left(z\in\pi_{0}\,y\right)\times\left(\operatorname{Acc}z\right)}\pi_{0}\circ e\sim\pi_{0}$ (76)
$\simeq\sum_{e\,:\,\sum_{z:V_{\operatorname{Acc}}}\pi_{0}\,z\in\pi_{0}\,x\ \simeq\ \sum_{z:V_{\operatorname{Acc}}}\pi_{0}\,z\in\pi_{0}\,y}\pi_{0}\circ e\sim\pi_{0}$ (77)
$\simeq\prod_{z:V_{\operatorname{Acc}}}z\in_{\operatorname{Acc}}x\simeq z\in_{\operatorname{Acc}}y$ (78)
In step (73) we use Lemma 6. Step (74) is extensionality for $(V,\in)$ and step (75) is Corollary 2. In step (76) we use the fact that elements of accessible sets are accessible and thus $\operatorname{Acc}z$ is contractible. Step (77) is some rearranging of conjuncts in the base type together with the characterisation of equality in subtypes for the fibration. The last step is again an application of Corollary 2.
Chasing $\operatorname{refl}$ along this chain of equivalences we see that it is sent to $\lambda\,z.\operatorname{id\mbox{-}equiv}$.∎
The subtype of accessible elements satisfies foundation.
Theorem 2 ( ).
Given an $\in$-structure, $(V,\in)$, the $\in$-structure $(V_{\operatorname{Acc}},\in_{\operatorname{Acc}})$ has foundation.
Proof.
There are two different accessibility predicates at play here. Let $\operatorname{Acc}$ and $\operatorname{acc}$ be accessibility with respect to $(V,\in)$ and let $\operatorname{Acc}^{\prime}$ and $\operatorname{acc}^{\prime}$ be accessibility with respect to $(V_{\operatorname{Acc}},\in_{\operatorname{Acc}})$. We need to show $\prod_{x:V_{\operatorname{Acc}}}\operatorname{Acc}^{\prime}x$, which we do by using the induction principle of $\operatorname{Acc}$ (modulo a transport):
$\alpha:\prod_{x:V_{\operatorname{Acc}}}\operatorname{Acc}^{\prime}x$
$\alpha(x,\operatorname{acc}\,f):=\operatorname{acc}^{\prime}\,\left(\lambda\,(y:V_{\operatorname{Acc}})(p:\pi_{0}\,y\in x).\,\alpha(\pi_{0}\,y,f\,(\pi_{0}\,y)\,p)\right)\qed$
3 Internalisations of types in $\in$-structures
Any element $a:V$ gives rise to a type $\operatorname{El}a$ (Definition 2), and in some sense $a$ represents $\operatorname{El}a$ inside the bigger structure of $(V,\in)$. For instance, an operation from $a$ to $b$ is precisely a function from $\operatorname{El}a\rightarrow\operatorname{El}b$ (Proposition 7). A natural question to ask is: Which types can be represented as elements in this way in a given $\in$-structure? In this section we introduce some basic vocabulary for talking about this kind of representation. We apply this by giving a very flexible formulation of the axiom of infinity – which constructively is often formulated as the existence of a set collecting the natural numbers for some chosen encoding of these. The flexibility of our formulation is that it takes the encoding as a parameter, making very few assumptions about it. We prove that for $\in$-structures that satisfy replacement, the existence of a set of natural numbers is independent of encoding.
3.1 Internalisations and representations
For the rest of this section, fix an $\in$-structure $(V,\in)$ and $A:\operatorname{Type}$.
Definition 16 ( ).
A $(V,\in)$-internalisation of $A$ is an element $a:V$ such that $\operatorname{El}a\simeq A$.
There can be several different internalisations of $A$. However, if we fix an encoding of the elements of $A$ in $V$, there is at most one element in $V$ which is an internalisation of $A$ with respect to this encoding.
An encoding of the elements of $A$ in $V$ is a function $A\rightarrow V$. We will call this a representation because of the superficial similarity with classical representation theory, which can be seen as the study of functors $G\rightarrow\operatorname{Vec}_{k}$ for some group $G$ and field $k$. In our case representations are functions $A\rightarrow V$ where $A$ can be any type (not necessarily a group), and the codomain is some $\in$-structure.
Definition 17 ( ).
A $(V,\in)$-representation of $A$ is a function $A\rightarrow V$.
Our main concern is going to be if a given representation gives rise to an internalisation of the domain in the $\in$-structure itself. In other words, if the elements pointed out by the representation can be collected to a set such that the type of elements is exactly the type being represented.
Definition 18 ( ).
Given a $(V,\in)$-representation of $A$, say $f:A\rightarrow V$, let an internalisation of $f$ be an $a:V$ such that for every $z:V$ we have $z\in a\simeq\operatorname{fiber}f\,z$.
We say that a representation is internalisable if there is an internalisation of it. An internalisation of a representation is an internalisation of the domain.
Proposition 10 ( ).
Let $f:A\rightarrow V$ be a $(V,\in)$-representation of $A$ and let $a:V$ be an internalisation of $f$. Then $a$ is an internalisation of $A$.
Proof.
By Lemma 4.8.2 in the HoTT Book [hottbook], we have:
$\operatorname{El}a\equiv\left(\sum_{z:V}z\in a\right)\simeq\left(\sum_{z:V}\operatorname{fiber}f\,z\right)\simeq A,$
showing that $a$ is an internalisation of $A$.∎
Once a representation of a type is fixed, the internalisation is uniquely defined. Thus, we will speak of the internalisation of a given representation.
Proposition 11 ( ).
Given a $(V,\in)$-representation of $A$, say $f:A\rightarrow V$, the internalisation of $f$ is uniquely defined. That is, the type $\sum_{a:V}\prod_{z:V}\left(z\in a\simeq\operatorname{fiber}f\,z\right)$ is a proposition.
Proof.
Corollary of Proposition 4.∎
The notion of representation gives us a way to separate the internalisation of a type into a structure part (how the elements are encoded) and the property that this structure can be internalised. In fact, an internalisation of a type is exactly the same as an internalisation of a representation of that type.
Proposition 12 ( ).
The type of all internalisable $(V,\in)$-representations of $A$, $\sum_{f:A\rightarrow V}\sum_{a:V}\prod_{z:V}z\in a\simeq\operatorname{fiber}f\,z$ is equivalent to the type of all internalisations of $A$, $\sum_{a:V}\left(\operatorname{El}\,a\simeq A\right)$.
Proof.
For $a:A$, we have the following chain of equivalences:
$\left(\sum_{f:A\rightarrow V}\prod_{z:V}z\in a\simeq\operatorname{fiber}f\,z\right)$ $\simeq\sum_{f:A\rightarrow V}\sum_{e:\operatorname{El}a\simeq A}f\circ e=\pi_{0}$ (79)
$\simeq\sum_{e:\operatorname{El}a\simeq A}\sum_{f:A\rightarrow V}f=\pi_{0}\circ e^{-1}$ (80)
$\simeq\left(\operatorname{El}a\simeq A\right)$ (81)
Step (79) is an application of Corollary 2. We rearrange in step (80), and finally, step (81) is the fact that the total space of paths from a given point is contractible.
The desired equivalence follows from the equivalence above, after some rearranging.∎
If a representation $f:A\rightarrow V$ is an embedding, we say that the representation is faithful. The truncation level of a representation determines the level of the internalisation. In the special case of a faithful representation, the internalisation is a mere set.
Proposition 13 ( ).
An internalisation of $f:A\rightarrow V$ is a $(k+1)$-type if and only if $f$ is a $k$-truncated representation, for $k:\mathbb{N}_{-1}$.
Proof.
Let $a:V$ be an internalisation of $f$. Then $a$ is a $(k+1)$-type if and only if $\operatorname{fiber}f\,z$ is a $k$-type, for all $z:V$. In other words, $a$ is a $(k+1)$-type if and only if $f$ is a $k$-truncated map.∎
Note here that if $f$ is a faithful representation, it is not necessarily the case that the type $A$ itself is a set. A mere set in $V$ can be the internalisation of a faithful representation of a type of any level, as long as it can be embedded into $V$. An exception to this is if $V$ is of level 0, meaning that it only contains mere sets, then $V$ itself is a set, all internalisable representations are faithful, and hence any type with an internalisable representation will be a set.
Representations and internalisations give an new perspective on the properties of Section 2. All of them, except foundation, can be seen as stating that a certain representation has an internalisation. For instance, exponentiation can be seen as stating that the representation sending a map $\operatorname{El}\,a\rightarrow\operatorname{El}\,b$ to its graph, can be internalised.
Replacement, in particular, can be seen in an interesting light. In classical set theory, replacement is the axiom schema which says that if you have a set you can replace its elements and still get a set, as long as the replacements are uniquely defined (by a formula in the first-order language of set theory). In terms of representations we can view this as follows: Given an internalisation of a type as a set, any other representation can be internalised by replacing the elements of the internalisation by their alternate representation. The various levels of replacement properties correspond to restrictions on what kind of representations can be replaced. For instance, 0-replacement gives the replacement property for faithful representations.
Proposition 14 ( ).
If $(V,\in)$ satisfies $(k+1)$-replacement and $A$ has an internalisation, then any $k$-truncated representation $f:A\rightarrow V$ of $A$ has an internalisation.
Proof.
Let $a:V$ be an internalisation of $A$, with a given $\alpha:\operatorname{El}a\simeq A$. Then apply $(k+1)$-replacement to $f\circ\alpha:\operatorname{El}a\rightarrow V$ to obtain $b:V$ such that for all $z:V$ there is an equivalence
$z\in b\simeq\Big{\|}\sum_{x:\operatorname{El}\,a}f(\alpha\,x)=z\Big{\|}_{k}$ (82)
Since $f$ is $k$-truncated and $\alpha$ is an equivalence, we can swap the base type $\operatorname{El}\,a$ with $A$ and drop the truncation, giving us an equivalence
$z\in b\simeq\sum_{x:A}f\,x=z$ (83)
Thus $b$ internalises $f$.∎
3.2 Natural Numbers / Infinity
There are many ways of formulating the axiom of infinity. The simplest formulation is perhaps $\exists\,u\,(\emptyset\in u\wedge\forall\,x\in u\ \exists\,y\in u\ x\in y)$, which can be found in some texts, such as Set Theory by Bell [bell2011]. This formulation depends on foundation for the result to actually be an infinite set, as this axiom is also satisfied by the co-hereditarily finite sets (the final coalgebra of the the finite powerset functor), by the element defined by the equation $x=\{\emptyset,x\}$. Another approach is to define the notion of a successor set, for instance $s\,x=x\cup\{x\}$ or $s\,x=\{x\}$ and then postulate the existence of a set containing $\emptyset$ and which is closed under $s$. This becomes an infinite set even without foundation, since the successor function preserves well-foundedness and increases rank.
In Aczel’s CZF [aczel1978], the axiom of infinity is formulated as the existence of a set of natural numbers, $\exists\,z\,\operatorname{Nat}(z)$. Aczel’s formulation of this axiom determines the set of natural numbers uniquely if foundation is assumed. However, if foundation is not assumed, one may have fixed points of the successor function which can be thrown in while still satisfying Aczel’s $\operatorname{Nat}$ predicate. For instance, given a quine atom, $q=\{q\}$, then for any $n$ we have $\operatorname{Nat}(n)\rightarrow\operatorname{Nat}(n\cup q)$. This can be remedied by further assuming that $n$ is accessible, in which case $n$ is uniquely determined.
When choosing a property corresponding to the axiom infinity for $\in$-structures in general, we will leverage the fact that we are not bound by first-order logic, and try to give a direct and intuitive formulation, which does not depend on foundation or assumptions about accessibility. We will also keep to the principle that properties postulating existence of sets should be uniquely determined, or be explicitly given as extra structure. Therefore, we will not say that an $\in$-structure has natural numbers to mean that $\mathbb{N}$ has an internalisation, as the type $\sum_{a:V}\operatorname{El}a\simeq\mathbb{N}$ is not a proposition. Having an internalisation of a type is a structure, rather than a property. However, having an internalisation with respect to a fixed representation is a property (Proposition 11), and it is the one we will use.
We will follow Aczel in choosing the natural numbers as the canonical infinite set, but leave the exact encoding of the natural numbers as extra structure. This leaves room for some exotic representations of the natural numbers, as the internalisation might not even be a mere set. Once an encoding is given, the set of naturals is uniquely determined, and will from a certain point of view behave like the usual natural numbers.
Definition 19 ( ).
Given an $\in$-structure $(V,\in)$ with a representation of $\mathbb{N}$, $f:\mathbb{N}\rightarrow V$, we say that $(V,\in)$ has natural numbers represented by $f$ if $f$ has an internalisation.
Suppose $n$ is a set of natural numbers represented by $f$ and let $e:\operatorname{El}n\simeq\mathbb{N}$ be the equivalence given by Proposition 10. Then we can define zero as
$\operatorname{zero}:\operatorname{El}n$ (84)
[…]
and we can define the successor function as
$\operatorname{suc}:\operatorname{El}n\rightarrow\operatorname{El}n$
$\operatorname{suc}x=e^{-1}\circ\operatorname{s}\circ e$
where $\operatorname{s}$ is the successor function on $\mathbb{N}$.
The usual induction principle holds, with respect to $\operatorname{zero}$ and $\operatorname{suc}$.
Proposition 15 ( ).
Given an $\in$-structure $(V,\in)$ with natural numbers $n:V$ represented by $f:\mathbb{N}\rightarrow V$, let $P:\operatorname{El}n\rightarrow\operatorname{Type}$ be a type family on $\operatorname{El}n$. Given $P\,\operatorname{zero}$ and $\prod_{x:\operatorname{El}n}P\,x\rightarrow P\,(\operatorname{suc}\,x)$, there is an element of the type $\prod_{x:\operatorname{El}n}P\,x$.
Proof.
Let $e:\operatorname{El}n\simeq\mathbb{N}$ be the equivalence given by Proposition 10. The result follows from the induction principle on $\mathbb{N}$, transported along $e$.∎
Example 1 ( ).
[…]
Example 2.
[…]
Interestingly, as long as there is enough replacement in the $\in$-structure, the exact choice of representation of $\mathbb{N}$ does not matter:
Proposition 16 ( ).
If $(V,\in)$ satisfies $(k+1)$-replacement and has natural numbers for some representation, then $(V,\in)$ has natural numbers represented by $f$ for any $k$-truncated representation $f$.
Proof.
Suppose $(V,\in)$ has natural numbers for some representation. By Proposition 10 it follows that $\mathbb{N}$ has an internalisation. Thus the result follows by Proposition 14.∎
4 Equivalence of extensional coalgebras and U-like $\in$-structures
There is a well-established coalgebraic reading of set theory in which models of set theory can be understood as coalgebras for the powerset functor on the category of classes. This can be traced back to Rieger [rieger1957], but perhaps found more explicitly in work by Osius [osius1974].§§§More recent work in this direction is the work of Paul Taylor [taylor2023] studying coalgebraic notions of well-foundedness and recursion. In this section we establish a similar correspondence between $\in$-structures and coalgebras. We define a hierarchy of functors $\operatorname{P}^{n}_{U}$, relative to a universe, and stratified by type levels, whose extensional coalgebras correspond to $\in$-structures at the given type level.
Definition 20 ( ).
For $n:\mathbb{N}_{-1}^{\infty}$, define $\operatorname{P}_{U}^{n+1}:\operatorname{Type}\rightarrow\operatorname{Type}$ by $\operatorname{P}_{U}^{n+1}\,X=\sum_{A:U}A\hookrightarrow_{n}X$.
Remark: $\operatorname{P}^{\infty}_{U}$ is a polynomial functor, but for finite $n$, $\operatorname{P}^{n}_{U}$ is not polynomial. Note that $\operatorname{P}^{0}_{U}$ is a $U$-restricted ‘powerset functor’ in that $\operatorname{P}^{0}_{U}\,X$ is the type of $U$-small subtypes of $X$.
The functorial action of $\operatorname{P}^{n}_{U}$ on a function $f:X\rightarrow Y$ is to postcompose with $f$ and then take the $(n-1)$-image and the $(n-1)$-image inclusion. There is a size issue with this though. With the usual construction of the $n$-image as $\sum_{b:B}\|\sum_{a:A}g\,a=b\|_{n}$ [hottbook], this type is in $U$ if $U$ is closed under truncation and if both the domain and the codomain are in $U$. In our case the domain is in $U$, but the codomain need not be. Is it possible to weaken the assumption that the codomain is in $U$ (given that the domain is) and still conclude that the $n$-image is in $U$?
4.1 Images of small types
As a higher groupoid, the image of a function $f:X\rightarrow Y$ looks like it has the points from $X$, but the paths are the paths from $Y$ resulting from applying $f$. For example, consider the function $f:\mathbf{1}\rightarrow\operatorname{Type}$ given by $f\,x:=\mathbf{2}$. The image of $f$ is the type $\sum_{x:\operatorname{Type}}\|\mathbf{2}=x\|_{-1}$. There is one point, namely $(\mathbf{2},|\operatorname{refl}|_{-1})$. However, by univalence, there are two distinct paths $(\mathbf{2},|\operatorname{refl}|_{-1})=(\mathbf{2},|\operatorname{refl}|_{-1})$: the identity equivalence and the equivalence that flips the elements in $\mathbf{2}$ (we only need to consider equality on the first coordinate).
More generally, the $n$-image looks like it has the $k$-cells, for $k\leq n+1$, from $X$, but the $(n+2)$-cells are the $(n+2)$-cells from $Y$ between the $(n+1)$-cells resulting from applying $f$. From this intuition it seems reasonable that if all the $(n+2)$-iterated identity types of $Y$ lie in $U$, along with $X$, then the resulting higher groupoid $n$-image also lies in $U$.
It turns out that it is enough to assume that this holds for $(-1)$-images. As pointed out below, this is not true for all univalent universes, but the results presented here will rely upon the assumption that this holds for our particular universe.
Rijke’s modified join construction [rijke2017] proves that a sufficient criterion for this smallness of images assumption to hold in a univalent universe is that $U$ is closed under homotopy colimits (or has graph quotients), and that global function extensionality holds. Here, we have chosen instead to directly assume that images of $U$-small types into locally $U$-small types are $U$-small.
The formulation of this assumption relies on two natural notions of smallness — which are also found in Rijke’s work [rijke2017] along with the fact that being small a proposition. We generalise to a definition of local smallness at any level.
Definition 21 ( ).
A type $X$ is essentially $U$-small if there is $A:U$ such that $A\simeq X$. That is, $\sum_{A:U}A\simeq X$.
We proceed by induction to define local smallness at higher levels.
Definition 22 ( ).
Let $X$ be a type. We define the notion of $X$ being $n$-locally $U$-small, for $n:\mathbb{N}^{\infty}$, as follows:
  • $X$ is 00-locally $U$-small if it is essentially $U$-small.
  • $X$ is $(n+1)$-locally $U$-small if for all $x,y:X$, the identity type $x=y$ is $n$-locally $U$-small.
  • $X$ is $\infty$-locally $U$-small.
We say that a $1$-locally $U$-small type is locally $U$-small.
Lemma 7 ( ).
Being essentially $U$-small is a mere proposition.
Proof.
We must show that the type $\sum_{A:U}A\simeq X$ is a mere proposition. This type looks close to being of the form $\sum_{b:T}b=a$, which is known to be contractible, but since $X$ is not a type in the universe we cannot directly apply univalence. Instead we give a direct proof based on the definition of being a mere proposition.
Let $(A,\alpha)$ and $(B,\beta)$ be two elements of $\sum_{A:U}A\simeq X$. Applying univalence, we observe that $\operatorname{ua}\,(\beta^{-1}\cdot\alpha):A=B$.
It remains to show that we get $\beta$ by transporting $\alpha$ along the path $\operatorname{ua}\,(\beta^{-1}\cdot\alpha)$, in the family $\lambda(Y:U).Y\simeq X$. But this is easily computed with path algebra:
$\operatorname{tr}^{\lambda Y.Y\simeq X}_{\operatorname{ua}\,(\beta^{-1}\cdot\alpha)}\alpha$ $=\alpha\cdot(\beta^{-1}\cdot\alpha)^{-1}$
$=\alpha\cdot\alpha^{-1}\cdot\beta$
$=\beta\qed$
Now, the following assumption, which will be assumed throughout the rest of the paper, can be formulated:
Assumption 1 (Images of small types ).
For every small type $A:U$ and every locally $U$-small $X:Type$, and given a function $f:A\rightarrow X$, there is a $U$-small type $\operatorname{image}\,f:U$, and functions $\operatorname{surj}f:A\twoheadrightarrow\operatorname{image}\,f$ and $\operatorname{incl}f:\operatorname{image}\,f\hookrightarrow X$, such that for every $a:A$ we have $\operatorname{incl}f\ (\operatorname{surj}f\,x)\equiv f\,x$.
Remark: Not all univalent universes satisfy the assumption above. For instance, given any univalent universe, $U$, we can define the subuniverse of mere sets, $\operatorname{Set}_{U}$. This subuniverse is univalent, and locally $\operatorname{Set}_{U}$-small, but the image of the map $(\lambda\,x.\mathbf{2}):\mathbf{1}\rightarrow\operatorname{Set}_{U}$ is not a set, as we saw above. Thus, the image of this map is not essentially $\operatorname{Set}_{U}$-small.
While the work in this article is presented informally in type theory, there might be some benefit here to spell out the assumption as formulated in Agda. This formulation is uniform in universe levels (the parameters i and j) so that it can be applied at any level.
module _ {i j}
         {Domain : Type i} {Codomain : Type j}
         (_ : is-locally-small i Codomain)
         (f : Domain  Codomain) where
  postulate Image : Type i
  postulate image-inclusion : Image  Codomain
  postulate image-quotient : Domain  Image
  postulate image-
     :  x  (image-inclusion  image-quotient  x  )  f x
  {-# REWRITE image- #-}
This code can be found in the project source repository, where the file containing the above snippet is called image-factorisation.agda ( ).
From the assumption of small $(-1)$-images, the smallness of $n$-images follows.
Proposition 17 ( ).
Let $n:\mathbb{N}^{\infty}_{-2}$. For every small type $A:U$ and every $(n+2)$-locally $U$-small $X:\operatorname{Type}$, and given a function $f:A\rightarrow X$, there is a small type $\operatorname{image}_{n}\,f:U$, together with an $n$-connected map $\operatorname{surj}_{n}f:A\twoheadrightarrow_{n}\operatorname{image}_{n}\,f$ and an $n$-truncated map $\operatorname{incl}_{n}f:\operatorname{image}_{n}\,f\hookrightarrow_{n}X$, such that for every $a:A$ we have $\operatorname{incl}f\ (\operatorname{surj}f\,x)=f\,x$.
Proof.
This is Proposition 2.2 in [christensen_non-accessible_2021], but we will also give our own proof here, formulated in a slightly different way. Let $A:U$ and $X:Type$ be $n$-locally $U$-small, and let $f:A\rightarrow X$. We proceed by induction on $n$.
For the base case we need to construct a type $\operatorname{image}_{-2}\,f:U$ together with a $(-2)$-connected map $\operatorname{surj}_{-2}f:A\twoheadrightarrow_{-2}\operatorname{image}\,f$ and a $(-2)$-truncated map $\operatorname{incl}_{-2}f:\operatorname{image}\,f\hookrightarrow_{-2}X$. Note that any map is $(-2)$-connected and that $(-2)$-truncated maps are precisely equivalences. By assumption $X$ is essentially $U$-small, i.e. there is a type $X^{\prime}:U$ together with an equivalence $e:X\simeq X^{\prime}$. So we take $\operatorname{image}_{-2}\,f:=X^{\prime}$, $\operatorname{surj}_{-2}f:=e\circ f$ and $\operatorname{incl}_{-2}f:=e^{-1}$.
Suppose the proposition holds for $n$, and suppose $X$ is $(n+1)$-locally $U$-small. Let $\operatorname{image}^{HoTT}_{n}g$ denote the HoTT Book definition of the $n$-image of a function $g$ [hottbook, Def. 7.6.3]. We will show that $\operatorname{image}^{HoTT}_{n-1}f$ is essentially $U$-small. By Theorem 7.6.6 in [hottbook] we have a factorisation
${A}$${X}$${\operatorname{image}^{HoTT}_{n-1}f}$$\scriptstyle{f}$$\scriptstyle{s}$$\scriptstyle{i}$
where $s$ is $(n-1)$-connected and $i$ is $(n-1)$-truncated. The map $s$ is in particular surjective, so by the uniqueness of the image factorisation, $\operatorname{image}^{HoTT}_{n-1}f\simeq\operatorname{image}\,s$. Thus we are done if we can show that $\operatorname{image}^{HoTT}_{n-1}f$ is locally $U$-small. Since being $U$-small is a proposition and $s$ is surjective, it is enough to show that $s\,x=s\,y$ is essentially $U$-small for all $x,y:A$. But by the characterisation of equality in $n$-images we have the equivalence $\left(s\,x=s\,y\right)\simeq\operatorname{image}^{HoTT}_{n-2}\,\left(\operatorname{ap}_{f}\,x\,y\right)$. The domain $x=y$ of $\operatorname{ap}_{f}\,x\,y$ is in $U$ and the codomain $f\,x=f\,y$ is $n$-locally $U$-small by assumption. Thus, by the induction hypothesis and the uniqueness of the $(n-2)$-image factorisation, $\operatorname{image}^{HoTT}_{n-2}\,\left(\operatorname{ap}_{f}\,x\,y\right)\simeq\operatorname{image}_{n-2}\,\left(\operatorname{ap}_{f}\,x\,y\right)$ and thus $s\,x=s\,y$ is essentially $U$-small.
For the case $n=\infty$, note that an $\infty$-connected map is an equivalence and that any map is $\infty$-truncated. Thus we take $\operatorname{image}_{\infty}\,f:=A$, $\operatorname{surj}_{\infty}\,f:=\operatorname{id\mbox{-}equiv}$ and $\operatorname{incl}_{\infty}\,f:=f$.∎
4.2 The functor $\operatorname{P}^{n}_{U}$
With the construction of $U$-small $n$-images, we can define the functorial action of $\operatorname{P}^{n}_{U}$. For $n:\mathbb{N}^{\infty}$, and two types $X,Y$ such that $Y$ is $(n+1)$-locally $U$-small, the $(n-1)$-image of any function $\phi:X\rightarrow Y$, $\operatorname{image}_{n-1}\,\phi$, lies in $U$ by Proposition 17. Moreover, the image inclusion $\operatorname{incl}_{n-1}\,\phi:\operatorname{image}_{n-1}\,\phi\rightarrow Y$ is an $(n-1)$-truncated map. Thus we make the following definition:
Definition 23 ( ).
Let $n:\mathbb{N}^{\infty}$. For types $X,Y$ such that $Y$ is $(n+1)$-locally $U$-small, and a function $\phi:X\rightarrow Y$ let $\operatorname{P}^{n}_{U}\,\phi:\operatorname{P}^{n}_{U}\,X\rightarrow\operatorname{P}^{n}_{U}\,Y$ be the function
$\operatorname{P}^{n}_{U}\phi\ (A,f):=\left(\operatorname{image}_{n-1}\,(\phi\circ f),\operatorname{incl}_{n-1}\,(\phi\circ f)\right)$
Note: For $n=\infty$ we get $\operatorname{P}^{\infty}_{U}\phi\ (A,f)\equiv(A,\phi\circ f)$.
Thinking of the elements in $\operatorname{P}^{n}_{U}\,X$ as slices over the type $X$, equality should be fiberwise equivalence. This is indeed the case, as the next proposition shows.
Proposition 18 ( ).
Given $n:\mathbb{N}^{\infty}$ and a type $X$, for $(A,f),(B,g):\operatorname{P}^{n}_{U}\,X$ there is an equivalence
$\left((A,f)=(B,g)\right)\simeq\prod_{x:X}\operatorname{fiber}\,f\,x\simeq\operatorname{fiber}\,g\,x$
Proof.
For $(A,f),(B,g):\operatorname{P}^{\infty}_{U}\,X$ we have the following chain of equivalences:
$\left((A,f)=(B,g)\right)$ $\simeq\sum_{\alpha:A=B}g\circ(\operatorname{coe}\,\alpha)=f$ (86)
$\simeq\sum_{\alpha:A\simeq B}g\circ\alpha=f$ (87)
$\simeq\sum_{\alpha:\sum_{x:X}\operatorname{fiber}\,f\,x\;\simeq\;\sum_{x:X}\operatorname{fiber}\,g\,x}\pi_{0}\circ\alpha=\pi_{0}$ (88)
$\simeq\prod_{x:X}\operatorname{fiber}\,f\,x\simeq\operatorname{fiber}\,g\,x$ (89)
Step (86) is the usual characterisation of equality in $\phi$-types, step (87) is the univalence axiom, step (88) is Lemma 4.8.2 in the HoTT Book [hottbook, p. 186] and step (89) is Corollary 2. (Note: the equivalence $\left(\sum_{\alpha:A\simeq B}g\circ\alpha=f\right)\simeq\prod_{x:X}(\operatorname{fiber}\,f\,x\simeq\operatorname{fiber}\,g\,x)$ was proved as Lemma 5 in Multisets in type theory [gylterud-multisets], by explicitly constructing an equivalence.)
For finite $n$, the type $\operatorname{P}^{n}_{U}\,X$ is a subtype of $\operatorname{P}^{\infty}_{U}\,X$, and thus has the same identity types.∎
The characterisation of equality on $\operatorname{P}^{n}_{U}$ allows us to compute the type level.
Corollary 3 ( ).
Given $n:\mathbb{N}^{\infty}$ and a type $X$, the type $\operatorname{P}^{n}_{U}\,X$ is an $n$-type.
Proof.
For $(A,f),(B,g):\operatorname{P}^{n}_{U}\,X$ and $x:X$, the types $\operatorname{fiber}\,f\,x$ and $\operatorname{fiber}\,g\,x$ are $(n-1)$-truncated. Thus the identity type $(A,f)=(B,g)$ is $(n-1)$-truncated.∎
The functor $\operatorname{P}^{n}_{U}$ preserves homotopies.
Proposition 19 ( ).
Let $n:\mathbb{N}^{\infty}$. For types $X,Y$ such that $Y$ is $(n+1)$-locally $U$-small, and functions $\phi,\psi:X\rightarrow Y$, any homotopy $\epsilon:\phi\sim\psi$ induces a homotopy $\operatorname{P}^{n}_{U}\,\epsilon:\operatorname{P}^{n}_{U}\phi\sim\operatorname{P}^{n}_{U}\psi$. Moreover, $\operatorname{refl\mbox{-}htpy}$ is sent to $\operatorname{refl\mbox{-}htpy}$.
Proof.
Given $\epsilon:\phi\sim\psi$ we construct the homotopy
$\operatorname{P}^{n}_{U}\epsilon:\operatorname{P}^{n}_{U}\phi\sim\operatorname{P}^{n}_{U}\psi$
$\operatorname{P}^{n}_{U}\epsilon\,(A,f):=\operatorname{ap}_{\lambda\,\sigma.\operatorname{P}^{n}_{U}\sigma\,(A,f)}(\operatorname{funext}\,\epsilon)$
For $\operatorname{P}^{n}_{U}\operatorname{refl\mbox{-}htpy}$ we have the following chain of equalities:
$\operatorname{P}^{n}_{U}\operatorname{refl\mbox{-}htpy}\,(A,f)$ $\equiv\operatorname{ap}_{\lambda\,\sigma.\operatorname{P}^{n}_{U}\sigma\,(A,f)}(\operatorname{funext}\,\operatorname{refl\mbox{-}htpy})$
$=\operatorname{ap}_{\lambda\,\sigma.\operatorname{P}^{n}_{U}\sigma\,(A,f)}\operatorname{refl}$
$\equiv\operatorname{refl}\qed$
We recap the definition of an algebra for the functor $\operatorname{P}^{n}_{U}$ and a homomorphism between two algebras.
Definition 24 ( ).
A $\operatorname{P}^{n}_{U}$-algebra, $(X,m)$ consists of an $(n+1)$-locally $U$-small type $X$ and a map $m:\operatorname{P}^{n}_{U}X\rightarrow X$.
The restriction to $(n+1)$-locally $U$-small carrier types for $\operatorname{P}^{n}_{U}$ algebras, is because $\operatorname{P}^{n}_{U}$ is only functorial on such types.
Definition 25 ( ).
Given two $\operatorname{P}^{n}_{U}$ algebras, $(X,m)$ and $(X^{\prime},m^{\prime})$, a $\operatorname{P}^{n}_{U}$-algebra homomorphism from $(X,m)$ to $(X^{\prime},m^{\prime})$ is a pair $(\phi,\alpha)$ where $\phi:X\rightarrow X^{\prime}$ and $\alpha:\phi\circ m\sim m^{\prime}\circ\operatorname{P}^{n}_{U}\phi$.
For two $\operatorname{P}^{n}_{U}$-algebra homomorphisms $(\phi,\alpha)$ and $(\psi,\beta)$ to be equal, there needs to be a homotopy between the underlying maps $\phi$ and $\psi$. But there is also a coherence condition on $\alpha$ and $\beta$, which can be seen as saying that the two natural ways of proving that $m^{\prime}\,(\operatorname{P}^{n}_{U}\,\phi)=_{X}^{\prime}\psi\,(m\,a)$ are equal identifications in $X^{\prime}$. This holds trivially if $X^{\prime}$ is a mere set, and in that case only the homotopy between the maps is relevant.
Lemma 8 ( ).
For any two $\operatorname{P}^{n}_{U}$-algebra homomorphisms $(\phi,\alpha)$ and $(\psi,\beta)$, from $(X,m)$ to $(X^{\prime},m^{\prime})$, their identity type is characterised by:
$\left((\phi,\alpha)=(\psi,\beta)\right)\simeq\sum_{\epsilon:\phi\sim\psi}\prod_{a:\operatorname{P}^{n}_{U}X}\alpha\,a\cdot\operatorname{ap}_{m^{\prime}}(\operatorname{P}^{n}_{U}\,\epsilon\,a)=\epsilon\,(m\,a)\cdot\beta\,a$ (90)
Proof.
We use the structure identity principle. Given an algebra homomorphism $(\phi,\alpha)$, we need to define families $P:(X\rightarrow X^{\prime})\rightarrow\operatorname{Type}$ and $Q:\prod_{\psi:X\rightarrow X^{\prime}}\left(\psi\circ m\sim m^{\prime}\circ\operatorname{P}^{n}_{U}\psi\right)\rightarrow P\,\psi\rightarrow\operatorname{Type}$, which should characterise the identity type of the base type and of the fibration respectively. Then we need to construct elements $p:P\,\phi$ and $q:Q\,\phi\,\alpha\,p$. Finally, we need to show that for all $\psi:X\rightarrow X^{\prime}$, $\left(\phi=\psi\right)\simeq P\,\psi$ and that for all $\alpha^{\prime}:\phi\circ m\sim m^{\prime}\circ\operatorname{P}^{n}_{U}\phi$, $\left(\alpha=\alpha^{\prime}\right)\simeq Q\,\phi\,\alpha^{\prime}\,p$.
The family $P$ is defined as $P\,\psi:=\left(\phi\sim\psi\right)$, and $Q$ is defined as $Q\,\psi\,\beta\,p^{\prime}:=\prod_{a:\operatorname{P}^{n}_{U}X}\alpha\,a\cdot\operatorname{ap}_{m^{\prime}}(\operatorname{P}^{n}_{U}\,p^{\prime}\,a)=p^{\prime}\,(m\,a)\cdot\beta\,a$. For the element in $P\,\phi$ we take $p:=\operatorname{refl\mbox{-}htpy}$, and for $Q\,\phi\,\alpha\,p$ and $(A,f):\operatorname{P}^{n}_{U}X$ we have the following chain of equalities:
$\alpha\,(A,f)\cdot\operatorname{ap}_{m^{\prime}}(\operatorname{P}^{n}_{U}\,p\,(A,f))$ $=\alpha\,(A,f)\cdot\operatorname{ap}_{m^{\prime}}\operatorname{refl}$ (91)
$=\alpha\,(A,f)$ (92)
$=p\,(m\,(A,f))\cdot\alpha\,(A,f)$ (93)
In step (91) we use Proposition 19.
It remains to construct the equivalences. By function extensionality, $\left(\phi=\psi\right)\simeq\left(\phi\sim\psi\right)$. Moreover, for $\alpha^{\prime}:\phi\circ m\sim m^{\prime}\circ\operatorname{P}^{n}_{U}\phi$ we have the following chain of equivalences:
$\left(\alpha=\alpha^{\prime}\right)$ $\simeq\left(\alpha\sim\alpha^{\prime}\right)$ (94)
$\simeq\prod_{a:\operatorname{P}^{n}_{U}X}\alpha\,a\cdot\operatorname{ap}_{m^{\prime}}\operatorname{refl}=\operatorname{refl}\cdot\alpha^{\prime}\,a$ (95)
$\simeq\prod_{a:\operatorname{P}^{n}_{U}X}\alpha\,a\cdot\operatorname{ap}_{m^{\prime}}(\operatorname{P}^{n}_{U}\,p\,a)=p\,(m\,a)\cdot\alpha^{\prime}\,a$ (96)
$\equiv Q\,\phi\,\alpha^{\prime}\,p$ (97)
Step (94) is function extensionality, and in step (96) we again use Proposition 19.∎
4.3 U-likeness
In classical set theory, the Mostowsky collapse relates any well-founded set-like relation to the well-founded hierarchy. A set-like relation is a class relation, say $R$, where $\{y\ |\ (x,y)\in R\}$ is a set for all $x$. The notion of “being small”, i.e. being a set, is also one of the basic distinctions in algebraic set theory [joyal_algebraic_1995].
In type theory, types take the role of classes while the universe $U$ provides a measure of smallness, akin to being a set in set theory. Hence, the notion of a $U$-like $\in$-structure will mirror the classical notion of an extensional, set-like relation.
Definition 26 ( ).
An $\in$-structure, $(V,\in)$ is $U$-like if $\operatorname{El}a$ is essentially $U$-small for every $a:V$.
Remark: To simplify notation, we will coerce $\operatorname{El}a:U$ in $U$-like $\in$-structures.
Proposition 20 ( ).
Being $U$-like is a mere proposition for any $\in$-structure, $(V,\in)$.
The interaction between internalisable types and $U$-likeness is straightforward:
Proposition 21 ( ).
An $\in$-structure, $(V,\in)$, is $U$-like if any only if for each internalisable $A:\operatorname{Type}$, the type $A$ is essentially $U$-small.
Proof.
We have the following chain of equivalences:
$\prod_{A:\operatorname{Type}}$ $\left(\sum_{a:V}\operatorname{El}a\simeq A\right)\rightarrow\sum_{X:U}A\simeq X$
$\simeq\prod_{a:V}\left(\sum_{A:\operatorname{Type}}\operatorname{El}a\simeq A\right)\rightarrow\sum_{X:U}\operatorname{El}a\simeq X$ (98)
$\simeq\prod_{a:V}\left(\sum_{X:U}\operatorname{El}a\simeq X\right)$ (99)
In step (99) we use the fact that $\sum_{A:\operatorname{Type}}\operatorname{El}a\simeq A$ is contractible.∎
We will now investigate in detail the relationship between $U$-like $\in$-structures and $\operatorname{P}^{n}_{U}$-coalgebras.
Lemma 9 ($\in$-structures are coalgebras ).
For a fixed $V$, having a $U$-like $\in$-structure on $V$ is equivalent to having a coalgebra structure $V\hookrightarrow\operatorname{P}^{\infty}_{U}\,V$, which is an embedding. This equivalence sends the relation $\in\,:V\rightarrow V\rightarrow\operatorname{Type}$ to the coalgebra $\lambda\,a.(\operatorname{El}\,a,\pi_{0})$ and, in the other direction, it sends the coalgebra $m:V\rightarrow\operatorname{P}^{\infty}_{U}\,V$ to the relation $\lambda\,b\,a.\operatorname{fiber}\ (\pi_{1}\,(m\,a))\ b$.
Proof.
We have the following chain of equivalences:
$V\rightarrow\sum_{A:U}(A\rightarrow V)$ $\simeq\left(V\rightarrow\sum_{F:V\rightarrow\operatorname{Type}}\sum_{E:U}\left(\left(\sum_{b:V}F\,b\right)\simeq E\right)\right)$ (100)
$\simeq\sum_{\in:V\rightarrow V\rightarrow\operatorname{Type}}\prod_{a:V}\sum_{E:U}\left(\left(\sum_{b:V}b\in a\right)\simeq E\right)$ (101)
The equivalence (100) follows by substituting $\sum_{A:U}(A\rightarrow V)$ with the equivalent type
$\sum_{F:V\rightarrow\operatorname{Type}}\sum_{E:U}(\sum_{b:V}F\,b)\simeq E$, since a fibration $A\rightarrow V$, with a total type $A:U$ is equivalent to a family over $V$ with $U$-small total type, as being $U$-small is a mere proposition. The equivalence (101) is essentially currying.
Chasing a coalgebra, respectively an $\in$-relation, along the chain of equivalences we see that it computes as stated.
The type $\sum_{\in:V\rightarrow V\rightarrow\operatorname{Type}}\prod_{a:V}\sum_{E:U}\left(\left(\sum_{b:V}b\in a\right)\simeq E\right)$ is exactly the type of $\in$-structures on $V$ which are $U$-like, except for the extensionality requirement. It remains to show that the coalgebra is an embedding if and only if the corresponding $\in$-relation, given by the equivalence above, is extensional.
Given a coalgebra $m:V\rightarrow\operatorname{P}^{\infty}_{U}\,V$, let $\in$ be the corresponding relation given by the equivalence. By Proposition 18 we have, for any $x,y:V$, an equivalence
$e:\left(m\,x=m\,y\right)\simeq\left(\prod_{z:V}z\in x\simeq z\in y\right)$
This equivalence sends $\operatorname{refl}$ to $\lambda\,z.\operatorname{id\mbox{-}equiv}$. Let $e^{\prime}:x=y\rightarrow\prod_{z:V}z\in x\simeq z\in y$ be the map defined by path induction as $e^{\prime}\,\operatorname{refl}:=\lambda\,z.\operatorname{id\mbox{-}equiv}$. The following diagram commutes:
${x=y}$${m\,x=m\,y}$${\prod_{z:V}z\in x\simeq z\in y}$$\scriptstyle{\operatorname{ap}_{m}}$$\scriptstyle{e^{\prime}}$$\scriptstyle{e}$
Since $e$ is an equivalence it follows that $\operatorname{ap}_{m}$ is an equivalence if and only if $e^{\prime}$ is an equivalence.∎
Lemma 10.
Given a $U$-like $\in$-structure, $(V,\in)$, and any $n:\mathbb{N}_{-1}$, the following are equivalent:
  1. 1.
    $b\in a$ is an $n$-type for all $a,b:V$,
  2. 2.
    $\lambda\,a.(\operatorname{El}\,a,\pi_{0})$ restricts to a map $V\hookrightarrow\operatorname{P}^{n+1}_{U}\,V$.
Proof.
The fibres of $\pi_{0}:(\sum_{b:V}b\in a)\rightarrow V$ are exactly the types $b\in a$.∎
The following theorem combines the two previous lemmas in order to characterise $U$-like $\in$-structures of various levels in terms of coalgebras.
Theorem 3 ( ).
For a fixed $V$ and for $n:\mathbb{N}^{\infty}_{-1}$, having a $U$-like $\in$-structure on $V$ such that $b\in a$ is an $n$-type for all $a,b:V$ is equivalent to having a coalgebra structure $V\hookrightarrow\operatorname{P}^{n+1}_{U}\,V$.
Proof.
Simple corollary of the previous two lemmas.∎
5 Fixed-point models
Rieger’s theorem is a result in set theory which states that any class with a set-like binary relation, which is a fixed-point of the related powerset functor, is a model of ZFC- (ZF without foundation) [rieger1957]. In this section we prove an analogous result, considering the family of functors $\operatorname{P}^{n}_{U}$ as higher level generalisations of the powerset functor. Any fixed-point for $\operatorname{P}^{n}_{U}$ is in particular a $\operatorname{P}^{n}_{U}$-coalgebra for which the coalgebra map is an embedding. Thus, by Theorem 3, such a fixed-point gives rise to a $U$-like $\in$-structure. This induced $\in$-structure will satisfy almost all the properties defined in Section 2, at some level, the only exception being foundation.
So we assume in this section that we are given a type $V$ and an equivalence $\sup:\operatorname{P}^{n}_{U}V\simeq V$. Let $\operatorname{desup}:V\simeq\operatorname{P}^{n}_{U}V$ be the inverse of $\sup$. We will specifically show that the induced $\in$-structure on $V$ has the following properties:
  • Empty set.
  • $U$-restricted $n$-separation.
  • If $V$ is $(n+1)$-locally $U$-small, it has $\infty$-unordered $I$-tupling for all $(n-1)$-truncated types $I:U$.
  • If $V$ is $(k+1)$-locally $U$-small, for some $k\leq n$ then it has:
    • $k$-unordered $I$-tupling for all $I:U$,
    • $k$-replacement,
    • $k$-union.
  • $V$ has exponentiation for all ordered pairing structures.
  • $V$ has natural numbers represented by $f$ for any $(n-1)$-truncated representation $f:\mathbb{N}\rightarrow V$.
Notation: Given $x:V$, we will use the notation $\overline{x}:U$ and $\widetilde{x}:\overline{x}\hookrightarrow_{n-1}V$ for the type and $(n-1)$-truncated map such that $x=\sup\,(\overline{x},\widetilde{x})$, i.e. $\overline{x}=\pi_{0}\,(\operatorname{desup}\,x)$ and $\widetilde{x}=\pi_{1}\,(\operatorname{desup}\,x)$.
Definition 27 ( ).
Let $(V,\in)$ be the $\in$-structure on $V$ given by Theorem 3. Note that for $x,y:V$,
$y\in x\equiv\operatorname{fiber}\,\widetilde{x}\,y$
Proposition 22 ( ).
For any $x:V$ we have that $\overline{x}\simeq\operatorname{El}\,x$.
Proof.
We have the following:
$\operatorname{El}\,x\equiv\left(\sum_{y:V}\operatorname{fiber}\,\widetilde{x}\,y\right)\simeq\overline{x}$ (102)
where the equivalence in the last step is Lemma 4.8.2 in the HoTT Book [hottbook, p. 142].∎
Proposition 23 ( ).
The type $V$ is an $n$-type.
Proof.
$V\simeq\operatorname{P}^{n}_{U}V$ and $\operatorname{P}^{n}_{U}\,V$ is an $n$-type by Corollary 3.∎
Many of the constructions of the properties defined in Section 2 require us to use $\sup$ on the $k$-image of some function, and its $k$-truncated inclusion map. For this, the $k$-image needs to be $U$-small. We therefore need to assume that $V$ is appropriately locally $U$-small in several of the following theorems.
5.1 Empty set
The empty set is a special case of unordered tupling but it does not require any assumption about local smallness. Therefore we note this special case separately.
Theorem 4 ( ).
The $\in$-structure $(V,\in)$ has empty set.
Proof.
The empty type $\mathbf{0}$ embeds into any type, in particular, it embeds into $V$. Let $f:\mathbf{0}\rightarrow V$ be the unique map from $\mathbf{0}$ to $V$. Since the map is an embedding, it is also an $(n-1)$-truncated map. The empty set is defined as
$\emptyset:=\sup\,(\mathbf{0},f)$
For $x:V$ we have the following chain of equivalences:
$x\in\emptyset\simeq\operatorname{fiber}\,f\,x\simeq\left(\sum_{s:\mathbf{0}}f\,s=x\right)\simeq\mathbf{0}\qed$
5.2 Restricted separation
Theorem 5 ( ).
The $\in$-structure $(V,\in)$ has $U$-restricted $n$-separation.
Proof.
For $x:V$ and $\Phi:\operatorname{El}x\rightarrow(n-1)\operatorname{\mbox{-}Type}_{U}$ we construct the term $\{x\mid\Phi\}$ as follows:
$\{x\mid\Phi\}:=\sup\,\left(\sum_{a:\overline{x}}\Phi\,(\widetilde{x}\,a,(a,\operatorname{refl})),\lambda(a,\_).\widetilde{x}\,a\right).$
The map $\lambda(a,\_).\widetilde{x}\,a$ is $(n-1)$ truncated as it is the composition of the $(n-1)$-truncated map $\widetilde{x}$ and the projection $\pi_{0}:\sum_{a:\overline{x}}\Phi\,(\widetilde{x}\,a,(a,\operatorname{refl}))\rightarrow\overline{x}$, which is $(n-1)$-truncated since $\Phi$ is a family of $(n-1)$-truncated types.
For $z:V$ we then get the following chain of equivalences:
$z\in\{x\mid\Phi\}$ $=\sum_{(a,\_):\sum_{a:\overline{x}}\Phi\,(\widetilde{x}\,a,(a,\operatorname{refl}))}\widetilde{x}\,a=z$
$\simeq\sum_{(a,\_):\sum_{a:\overline{x}}\widetilde{x}\,a=z}\Phi\,(\widetilde{x}\,a,(a,\operatorname{refl}))$
$\simeq\sum_{e:(z\in x)}\Phi\,(z,e)\qed$
5.3 Unordered tupling
For unordered tupling we can construct the $k$-truncated version if $V$ is $(k+1)$-locally $U$-small, and we can construct the $\infty$-truncated version if the indexing type has lower type level than $V$.
Theorem 6 ( ).
Let $k:\mathbb{N}^{\infty}$ be such that $k\leq n$. If $V$ is $(k+1)$-locally $U$-small then the $\in$-structure $(V,\in)$ has k-unordered I-tupling for all $I:U$.
Proof.
Let $I:U$ and $v:I\rightarrow V$, and construct the element $\{v\}_{k}:V$ by
$\{v\}_{k}:=\sup\,(\operatorname{image}_{k-1}\,v,\operatorname{incl}_{k-1}\,v)$
Since $k\leq n$, $\operatorname{incl}_{k-1}\,v$ is an $(n-1)$-truncated map. Then, for any $z:V$, we have the following chain of equivalences:
$z\in\{v\}_{k}$ $=\operatorname{fiber}\,(\operatorname{incl}_{k-1}\,v)\ z$
$\simeq\big{\|}\operatorname{fiber}\,v\ z\,\big{\|}_{k-1}$
$\equiv\big{\|}\sum_{i:I}v\,i=z\,\big{\|}_{k-1}\qed$
Corollary 4 ( ).
Let $k:\mathbb{N}^{\infty}$ be such that $k\leq n$. If $V$ is $(k+1)$-locally $U$-small then $V$ contains the $k$-unordered $1$-tupling $\{x\}_{k}$, and the $k$-unordered $2$-tupling $\{x,y\}_{k}$, for any $x,y:V$.
Theorem 7 ( ).
If $V$ is $(n+1)$-locally $U$-small then it has an ordered pairing structure.
Proof.
Follows from the previous corollary together with Theorem 4, Theorem 1 and Proposition 23.∎
Theorem 8 ( ).
Let $k:\mathbb{N}$ be such that $k<n$. If $V$ is $(n+1)$-locally $U$-small, then the $\in$-structure $(V,\in)$ has $\infty$-unordered I-tupling for all $I:k\operatorname{\mbox{-}Type}_{U}$.
Proof.
Let $I:U$ and $v:I\rightarrow V$. We construct the element $\{v\}_{\infty}:V$ as in Theorem 6:
$\{v\}_{\infty}:=\sup\,(\operatorname{image}_{n-1}\,v,\operatorname{incl}_{n-1}\,v)$
Then, for any $z:V$, we have the following chain of equivalences:
$z\in\{v\}_{\infty}$ $=\operatorname{fiber}\,(\operatorname{incl}_{k-1}\,v)\ z$ (103)
$\simeq\big{\|}\operatorname{fiber}\,v\ z\,\big{\|}_{n-1}$ (104)
$\equiv\big{\|}\sum_{i:I}v\,i=z\,\big{\|}_{n-1}$ (105)
$\simeq\sum_{i:I}v\,i=z$ (106)
where step (106) is the fact that $\sum_{i:I}v\,i=z$ is $(n-1)$-truncated since $V$ is an $n$-type and $I$ is a $k$-type for some $k<n$.∎
5.4 Replacement
Theorem 9 ( ).
Let $k:\mathbb{N}^{\infty}$ be such that $k\leq n$. If $V$ is $(k+1)$-locally $U$-small, then the $\in$-structure $(V,\in)$ has $k$-replacement.
Proof.
For $x:V$ and $f:\operatorname{El}\,x\rightarrow V$, let $\phi:\overline{x}\rightarrow V$ be the function that sends $a:\overline{x}$ to $f\,(a,\operatorname{refl})$. We define the following element:
$\{f(y)\mid y\in x\}:=\sup\,(\operatorname{image}_{k-1}\,\phi,\operatorname{incl}_{k-1}\,\phi)$
Since $k\leq n$, $\operatorname{incl}_{k-1}\,\phi$ is an $(n-1)$-truncated map. For $z:V$ we have the following chain of equivalences:
$z\in\{f(y)\mid y\in x\}$ $=\operatorname{fiber}\,(\operatorname{incl}_{k-1}\,\phi)\,z$ (107)
$\simeq\;\big{\|}\operatorname{fiber}\,\phi\ z\big{\|}_{k-1}$ (108)
$\simeq\;\big{\|}\sum_{s:\operatorname{El}\,x}f\,s=z\big{\|}_{k-1}$ (109)
where (109) uses Proposition 22.∎
5.5 Union
Theorem 10 ( ).
Let $k:\mathbb{N}^{\infty}$ be such that $k\leq n$. If $V$ is $(k+1)$-locally $U$-small, then the $\in$-structure $(V,\in)$ has $k$-union.
Proof.
For $x:V$, let $\phi_{x}:\sum_{a:\overline{x}}\overline{(\widetilde{x}\,a)}\rightarrow V$ be the function that sends $(a,b)$ to $\widetilde{(\widetilde{x}\,a)}\,b$. We then define the union by
$\bigcup_{k}x:=\sup\,(\operatorname{image}_{k-1}\,(\phi_{x}),\operatorname{incl}_{k-1}\,(\phi_{x}))$
Since $k\leq n$, $\operatorname{incl}_{k-1}\,(\phi_{x})$ is an $(n-1)$-truncated map. For $z:V$ we thus have the following chain of equivalences:
$z\in\bigcup_{k}x$ $=\operatorname{fiber}\,(\operatorname{incl}_{k-1}\,(\phi_{x}))\,z$ (110)
$\simeq\;\big{\|}\operatorname{fiber}\,\phi_{x}\,z\big{\|}_{k-1}$ (111)
$\simeq\;\big{\|}\sum_{a:\overline{x}}\;z\in\widetilde{x}\,a\big{\|}_{k-1}$ (112)
$\simeq\;\big{\|}\sum_{y:V}\;(z\in y)\times(y\in x)\big{\|}_{k-1},$ (113)
where (113) uses Proposition 22.∎
5.6 Exponentiation
The property of having exponentiation is relative to an ordered pairing structure. It turns out that the construction of ordered pairs does not matter. $(V,\in)$ has exponentiation for any ordered pairing structure. In order to prove this though, we first need some lemmas.
Lemma 11 ( ).
Let $A$ be a type and $B$ a type family over $A$. For any two functions $\phi,\psi:\prod_{a:A}B\,a$ there is an equivalence
$\left(\phi=\psi\right)\simeq\sum_{e:A\simeq A}\prod_{a:A}(a,\phi\,a)=(e\,a,\psi\,(e\,a))$
Moreover, this equivalence sends $\operatorname{refl}$ to $(\operatorname{id\mbox{-}equiv},\operatorname{refl\mbox{-}htpy})$.
Proof.
For $\phi,\psi:\prod_{a:A}B\,a$ we have the following chain of equivalences:
$\left(\phi=\psi\right)$ $\simeq\left(\phi\sim\psi\right)$ (114)
$\simeq\sum_{e:A\simeq A}\sum_{r:\operatorname{id}_{A}\sim e}\prod_{a:A}\operatorname{tr}^{B}_{r\,a}(\phi\,a)=\psi\,(e\,a)$ (115)
$\simeq\sum_{e:A\simeq A}\prod_{a:A}\sum_{p:a=e\,a}\operatorname{tr}^{B}_{p}(\phi\,a)=\psi\,(e\,a)$ (116)
$\simeq\sum_{e:A\simeq A}\prod_{a:A}(a,\phi\,a)=(e\,a,\psi\,(e\,a))$ (117)
Step (114) is function extensionality. In step (115) we use the fact that the type $\sum_{e:A\simeq A}\operatorname{id}_{A}\sim e$ is contractible, with $(\operatorname{id\mbox{-}equiv},\operatorname{refl\mbox{-}htpy})$ as the center of contraction. Finally, step (117) is the characterisation of equality in $\Sigma$-types.
Chasing $\operatorname{refl}$ along the chain of equivalences, we get:
$\operatorname{refl}$ $\mapsto\operatorname{refl\mbox{-}htpy}$
$\mapsto(\operatorname{id\mbox{-}equiv},\operatorname{refl\mbox{-}htpy},\operatorname{refl\mbox{-}htpy})$
$\mapsto(\operatorname{id\mbox{-}equiv},\lambda\,a.(\operatorname{refl},\operatorname{refl}))$
$\mapsto(\operatorname{id\mbox{-}equiv},\operatorname{refl\mbox{-}htpy})$
Lemma 12 ( ).
Let $\langle-,-\rangle:V\rightarrow V\rightarrow V$ be an ordered pairing structure on $V$. Given a small type $A:U$ and a type family $B:A\rightarrow\operatorname{Type}$ together with $(n-1)$-truncated maps $f:A\hookrightarrow_{n-1}V$ and $g:\prod_{a:A}B\,a\hookrightarrow_{n-1}V$, there is an $(n-1)$-truncated map
$\operatorname{graph}_{f,g}:\left(\prod_{a:A}B\,a\right)\hookrightarrow_{n-1}V$
Proof.
Given $\phi:\prod_{a:A}B\,a$, we first need to construct an element in $V$. To this end, we construct the $(n-1)$-truncated map
$F_{\phi}:A\hookrightarrow_{n-1}V$
$F_{\phi}\,a=\langle f\,a,g\,a\,(\phi\,a)\rangle$
This map is $(n-1)$-truncated as it is the composition of the maps $\langle$-,$\rangle$ - , - ⟩, $\lambda\,(a,b).(f\,a,g\,a\,b)$ and $\lambda\,a.(a,\phi\,a)$. The first is $(-1)$-truncated by assumption, and thus $(n-1)$-truncated. The second is $(n-1)$-truncated since $f$ is $(n-1)$-truncated and $g\,a$ is $(n-1)$-truncated for all $a:A$. To see that the last map is $(n-1)$-truncated, for any pair $(a,b):\sum_{a:A}B\,a$ we have the following chain of equivalences:
$\left(\sum_{a^{\prime}:A}(a^{\prime},\phi\,a^{\prime})=(a,b)\right)\simeq\left(\sum_{a^{\prime}:A}\sum_{p:a^{\prime}=a}\operatorname{tr}^{B}_{p}(\phi\,a^{\prime})=b\right)\simeq\left(\phi\,a=b\right)$
where we have used the fact that the type $\sum_{a^{\prime}:A}a^{\prime}=a$ is contractible, with $(a,\operatorname{refl})$ as the center of contraction. Since $V$ is $n$-truncated, it follows that $\sum_{a^{\prime}:A}(a^{\prime},\phi\,a^{\prime})=(a,b)$ is $(n-1)$-truncated, and thus $\lambda\,a.(a,\phi\,a)$ is an $(n-1)$-truncated map.
This gives us the underlying map
$\operatorname{graph}_{f,g}:\left(\prod_{a:A}B\,a\right)\rightarrow V$
$\operatorname{graph}_{f,g}\,\phi:=\sup\,(A,F_{\phi})$
What is left is to show that this map is $(n-1)$-truncated. For this we show that $\operatorname{ap}_{\operatorname{graph}_{f,g}}:\phi=\psi\rightarrow\operatorname{graph}_{f,g}\,\phi=\operatorname{graph}_{f,g}\,\psi$ is $(n-2)$-truncated for all $\phi,\psi:\prod_{a:A}B\,a$.
Let $\phi,\psi:\prod_{a:A}B\,a$, let $\alpha:\left(\phi=\psi\right)\simeq\sum_{e:A\simeq A}\prod_{a:A}(a,\phi\,a)=(e\,a,\psi\,(e\,a))$ be the equivalence given by Lemma 11 and let $\beta:\left(\sum_{e:A\simeq A}F_{\phi}\sim F_{\psi}\circ e\right)\simeq\left((A,F_{\phi})=(A,F_{\psi})\right)$ be the equivalence given by the usual characterisation of identity in $\Sigma$-types together with univalence and function extensionality. We start by observing that
$\operatorname{ap}_{\operatorname{graph}_{f,g}}\sim\operatorname{ap}_{\sup}\circ\ \beta\ \circ\ (\lambda\,(e,H).(e,\operatorname{ap}_{\lambda\,(a,b).\langle f\,a,g\,a\,b\rangle}\circ H))\circ\alpha$
All three of $\alpha$, $\beta$ and $\operatorname{ap}_{\sup}$ are equivalences, and hence $(n-2)$-truncated maps. For the last map, it is enough to show that for all $e:A\simeq A$ and all $a:A$
$\operatorname{ap}_{\lambda\,(a,b).\langle f\,a,g\,a\,b\rangle}:(a,\phi\,a)=(e\,a,\psi\,(e\,a))\rightarrow F_{\phi}\,a=F_{\psi}\,(e\,a)$
is $(n-2)$-truncated. But this follows from the fact that the map $\lambda\,(a,b).\langle f\,a,g\,a\,b\rangle$ is $(n-1)$-truncated, as it is the composition of two $(n-1)$-truncated maps.∎
Theorem 11 (Exponentiation ).
The $\in$-structure $(V,\in)$ has exponentiation, for any ordered pairing structure.
Proof.
Let $\langle-,-\rangle:V\rightarrow V\rightarrow V$ be an ordered pairing structure on $V$. Given $x,y:V$, we define the element
$y^{x}:V$
$y^{x}:=\sup\,(\overline{x}\rightarrow\overline{y},\operatorname{graph}_{\widetilde{x},\lambda\,\_.\widetilde{y}})$
For $f:V$ we then have the following chain of equivalences:
$f\in y^{x}$ $\simeq\sum_{\phi:\overline{x}\rightarrow\overline{y}}f=\operatorname{graph}_{\widetilde{x},\lambda\,\_.\widetilde{y}}\,\phi$ (118)
$\simeq\sum_{\phi:\overline{x}\rightarrow\overline{y}}\prod_{z:V}z\in f\simeq z\in\operatorname{graph}_{\widetilde{x},\lambda\,\_.\widetilde{y}}\,\phi$ (119)
$\simeq\sum_{\phi:\overline{x}\rightarrow\overline{y}}\prod_{z:V}z\in f\simeq\sum_{a:\overline{x}}\langle\widetilde{x}\,a,\widetilde{y}\,(\phi\,a)\rangle=z$ (120)
$\simeq\sum_{\phi:\operatorname{El}\,x\rightarrow\operatorname{El}\,y}\prod_{z:V}z\in f\simeq\sum_{a:\operatorname{El}\,x}\langle\pi_{0}\,a,\pi_{0}\,(\phi\,a)\rangle=z$ (121)
$\simeq\operatorname{operation}\,x\,y\,f$ (122)
In step (119) we use extensionality and in step (121) we use Proposition 22. Finally, in step (122) we use Proposition 6.∎
5.7 Natural numbers
Theorem 12 ( ).
The $\in$-structure $(V,\in)$ has natural numbers represented by $f$, for any $(n-1)$-truncated representation $f$.
Proof.
Let $f:\mathbb{N}\rightarrow V$ be an $(n-1)$-truncated representation of $\mathbb{N}$. We construct the following element:
$n:=\sup\,(\mathbb{N},f)$
Then for any $z:V$ we have
$z\in n\equiv\operatorname{fiber}\,\widetilde{n}\,z=\operatorname{fiber}\,f\,z\qed$
6 The initial $\operatorname{P}^{n}_{U}$-algebra
The most straightforward way to construct a fixed-point of a functor is by constructing its initial algebra. For polynomial functors, the initial algebras are well understood. They are called W-types and are used to create a variety of inductive data structures, such as models of set theory in type theory: The initial algebra for $\operatorname{P}^{\infty}_{U}$ formed the basis for Aczel’s setoid model of CZF in Martin-Löf type theory [aczel1978]. The initial algebras for $\operatorname{P}^{n}_{U}$, for finite $n$, are subtypes of this type.
However, the functor $\operatorname{P}^{n}_{U}$ is not a polynomial functor, for finite $n$. So, its initial algebra will not be a simple W-type. Looking carefully at the type $\sum_{A:U}A\hookrightarrow_{n-1}X$ one can see that it is not strictly positive, since the type $A\hookrightarrow_{n-1}X:=\sum_{f:A\rightarrow X}\prod_{x:X}\operatorname{is\mbox{-}}(n-1)\operatorname{\mbox{-}type\,}(\operatorname{fiber}\,f\,x)$ contains a negative occurrence of $X$. Being strictly positive is a usual requirement for inductive definitions, so it is a bit surprising that $\operatorname{P}^{n}_{U}$ still has an initial algebra.
In this section, we will construct the initial algebra for $\operatorname{P}^{n}_{U}$, which then becomes a fixed-point $\in$-structure of level $n$. One interesting aspect of doing this proof in HoTT is that the carrier for a $\operatorname{P}^{n}_{U}$-algebra can lie on any type level.
We start by recalling Aczel’s W-type.
Definition 28 ( ).
Let $V^{\infty}:=W_{A:U}A$, and denote its (uncurried) constructor $\sup^{\infty}:\operatorname{P}^{\infty}_{U}V^{\infty}\rightarrow V^{\infty}$.
Remark: The superscript of the name, $V^{\infty}$, indicates that there is no bound on the level of the type. In fact, $V^{\infty}$ has the same type level as $U$.
The pair $(V^{\infty},\sup^{\infty})$ is the initial algebra for the functor $\operatorname{P}^{\infty}_{U}:\operatorname{Type}\rightarrow\operatorname{Type}$, and hence a fixed-point of $\operatorname{P}^{\infty}_{U}$. Any fixed-point has a canonical coalgebra structure, and by Lemma 9 this gives rise to a $U$-like $\in$-structure. Thus, we will denote by $\in^{\infty}:V^{\infty}\rightarrow V^{\infty}\rightarrow\operatorname{Type}$, this elementhood relation on $V^{\infty}$. Note: by construction, $x\in^{\infty}\sup^{\infty}\,(A,f)\equiv\operatorname{fiber}\,f\,x$.
This is the $\in$-structure explored in Multisets in type theory [gylterud-multisets], which also introduced some of the axioms of the previous sections. In From multisets to sets in Homotopy Type Theory [gylterud-iterative] a 0-level $\in$-structure was carved out as a subtype of $V^{\infty}$. And we will now prove that this 0-level structure is, as previously suspected, the initial $\operatorname{P}^{0}_{U}$ algebra. In fact, we will generalise the construction to form a type $V^{n}$ for all $n:\mathbb{N}$ and show that $V^{n}$ is the initial algebra for $\operatorname{P}^{n}_{U}$.
6.1 Iterative $n$-types
The types $V^{n}$ form a stratified hierarchy of $\in$-structures, where $V^{\infty}$ is at the top. This is analogous to how the $n$-types in $U$ form a hierarchy with $U$ itself at the top.
Definition 29 ( ).
Given $n:\mathbb{N}_{-1}$, define by induction on $V^{\infty}$, a predicate $\operatorname{is\mbox{-}it\mbox{-}}n\operatorname{\mbox{-}type}:V^{\infty}\rightarrow\operatorname{Type}$ by:
$\operatorname{is\mbox{-}it\mbox{-}}n\operatorname{\mbox{-}type}\left(\sup^{\infty}\,(A,f)\right):=\left(\operatorname{is\mbox{-}}n\operatorname{\mbox{-}trunc\mbox{-}map\,}f\right)\times\left(\prod_{a:A}\operatorname{is\mbox{-}it\mbox{-}}n\operatorname{\mbox{-}type}(f\,a)\right)$
This predicate is propositional, and an element of $V^{\infty}$ for which the predicate is true is called an iterative $n$-type.
Remark: We could define iterative $(-2)$-types analogously, but there are no elements in $V^{\infty}$ which satisfy this predicate, as the existence of such an element would imply that $V^{\infty}$ is essentially $U$-small.
Definition 30 ( ).
For $n:\mathbb{N}_{-1}$, let $V^{n+1}:=\sum_{x:V^{\infty}}\operatorname{is\mbox{-}it\mbox{-}}n\operatorname{\mbox{-}type}x$ denote the type of iterative $n$-types, which is a subtype of $V^{\infty}$.
Even though the type $V^{n}$ is not itself $U$-small, it has $U$-small identity types
Proposition 24 ( ).
The type $V^{n}$ is locally $U$-small.
Proof.
First, we observe that, since being an iterative $(n-1)$-type is a proposition, $V^{n}$ is a subtype of $V^{\infty}$. It was shown in Multisets in type theory [gylterud-multisets, Lemma 3] that the latter is locally $U$-small. Since the equality of a subtype is the underlying equality of the base type, it follows that $V^{n}$ is locally $U$-small.∎
In the following proofs we will leave out elements of types which are propositions, unless they are necessary, in order to increase readability. For the full details of the proofs, please see the Agda formalisation.
Proposition 25 ( ).
The map $\sup^{\infty}:\operatorname{P}^{\infty}_{U}V^{\infty}\rightarrow V^{\infty}$ restricts to a map $\sup^{n}:\operatorname{P}^{n}_{U}V^{n}\rightarrow V^{n}$.
Proof.
Let $A:U$ and $f:A\hookrightarrow_{n-1}V^{n}$. The element $\sup^{\infty}\,(A,\pi_{0}\circ f)$ is an iterative $(n-1)$-type:
  • $\pi_{0}\circ f$ is $(n-1)$-truncated since $f$ is $(n-1)$-truncated, and $\pi_{0}$ is the inclusion of a subtype, which is an embedding, and thus $(n-1)$-truncated.
  • For every $a:A$ we have $\pi_{1}\,(f\,a):\operatorname{is\mbox{-}it\mbox{-}}(n-1)\operatorname{\mbox{-}type}(\pi_{0}\,(f\,a))$.
Thus we construct the map
$\operatorname{sup}^{n}:\operatorname{P}^{n}_{U}V^{n}\rightarrow V^{n}$
[…]
Proposition 26 ( ).
There is a map $\operatorname{desup}^{n}:V^{n}\rightarrow\operatorname{P}^{n}_{U}V^{n}$.
Proof.
Let $x:V^{n}$, without loss of generality we may assume that $x\equiv(\operatorname{sup}^{\infty}\,(A,f),(p,q))$ for some $A:U$, $f:A\rightarrow V^{\infty}$, $p:\operatorname{is\mbox{-}}(n-1)\operatorname{\mbox{-}trunc\mbox{-}map\,}f$ and $q:\prod_{a:A}\operatorname{is\mbox{-}it\mbox{-}}(n-1)\operatorname{\mbox{-}type}(f\,a)$. The following function:
$\lambda\,a.(f\,a,q\,a):A\rightarrow V^{n}$
is $(n-1)$-truncated. This is because the composition $\pi_{0}\circ(\lambda\,a.(f\,a,q\,a))$ is $(n-1)$-truncated, by $p$, and $\pi_{0}$ is $(n-1)$-truncated as it is the inclusion of a subtype. This implies that the right factor, $\lambda\,a.(f\,a,q\,a)$, is $(n-1)$-truncated. Thus we construct the map
$\operatorname{desup}^{n}:V^{n}\rightarrow\operatorname{P}^{n}_{U}V^{n}$
$\operatorname{desup}^{n}\,(\operatorname{sup}^{\infty}\,(A,f),(p,q)):=(A,\lambda\,a.(f\,a,q\,a))\qed$
Proposition 27 ( ).
The maps $\sup^{n}$ and $\operatorname{desup}^{n}$ form an equivalence.
Proof.
For $(\operatorname{sup}^{\infty}\,(A,f),(p,q)):V^{n}$ we have
$\pi_{0}\,(\operatorname{sup}^{n}\,$ $(\operatorname{desup}^{n}\,(\operatorname{sup}^{\infty}\,(A,f),(p,q))))$
$\equiv\pi_{0}\,(\operatorname{sup}^{n}\,(A,\lambda\,a.(f\,a,q\,a)))$
$\equiv(\operatorname{sup}^{\infty}\,(A,f))$
$\equiv\pi_{0}\,(\operatorname{sup}^{\infty}\,(A,f),(p,q))$
Thus, by the characterisation of equality in subtypes, $\sup^{n}\circ\operatorname{desup}^{n}\sim\operatorname{id}$.
For $A:U$ and $f:A\hookrightarrow_{n-1}V^{n}$ we have
$\pi_{0}\,(\operatorname{desup}^{n}\,(\operatorname{sup}^{n}\,(A,f)))\equiv A$
and
$\pi_{0}\,(\pi_{1}$ $(\operatorname{desup}^{n}\,(\operatorname{sup}^{n}\,(A,f))))$
$\equiv\pi_{0}\,(\pi_{1}(\operatorname{desup}^{n}\,(\operatorname{sup}^{\infty}\,(A,\pi_{0}\circ f),(\_,\pi_{1}\circ f))))$
$\equiv\lambda\,a.(\pi_{0}\,(f\,a),\pi_{1}\,(f\,a))$
$\equiv\pi_{0}\,f$
Thus by the characterisation of equality in $\Sigma$-types, and equality in subtypes, $\operatorname{desup}^{n}\circ\sup^{n}\sim\operatorname{id}$.∎
Theorem 13 ( ).
$V^{n}$ is a fixed-point to the functor $\operatorname{P}^{n}_{U}$.
Proof.
Corollary of the previous proposition.∎
Theorem 14.
There is an $\in$-structure $(V^{n},\in^{n})$, on $V^{n}$, of level $n$.
Proof.
This follows from Theorem 13 together with Theorem 3.∎
Remark: By construction, for $x,y:V^{n}$ we have $y\in^{n}x\equiv\pi_{0}\,y\in^{\infty}\pi_{0}\,x$.
Remark: The $\in$-structure $(V^{0},\in^{0})$ is exactly the $\in$-structure studied in From multisets to sets in Homotopy Type Theory [gylterud-iterative], which there was established to be equivalent to the model of set theory given in the HoTT Book [hottbook].
6.2 Induction principle and recursors for $V^{n}$
Vn is a composite type: a $\Sigma$-type over a W-type and a predicate involving induction on that W-type. Since it is not a primitive inductive type it does not come with a ready-to-use induction principle. But, to some extent, $\sup^{n}:\operatorname{P}^{n}_{U}\,V^{n}\rightarrow V^{n}$ acts as a constructor, for which we have a corresponding induction principle. The expected computation rules do not hold definitionally, but are instead proven to hold up to identity.
Proposition 28 (Elimination for Vn ).
Given any family $P:V^{n}\rightarrow\operatorname{Type}$ and a function
$\phi:\prod_{A:U}\prod_{f:A\hookrightarrow_{n-1}V^{n}}\left(\prod_{a:A}P\,(f\,a)\right)\rightarrow P\,(\sup^{n}\,(A,f))$
there is a function $\operatorname{elim}_{V^{n}}\,P\ \phi:\prod_{x:V^{n}}P\,x$.
Furthermore, there is a path $\operatorname{elim}_{V^{n}}\,P\ \phi\,(\sup^{n}(A,f))=\phi\,A\,f\,(\operatorname{elim}_{V^{n}}\,P\,\phi\circ f)$, for any $A:U$ and $f:A\hookrightarrow_{n-1}V^{n}$.
Proof.
Given $x:V^{n}$, we may assume that $x\equiv(\operatorname{sup}^{\infty}\,(A,f),(p,q))$ where $A:U$, $f:A\rightarrow V^{\infty}$, $p:\operatorname{is\mbox{-}}(n-1)\operatorname{\mbox{-}trunc\mbox{-}map\,}f$ and $q:\prod_{a:A}\operatorname{is\mbox{-}it\mbox{-}}(n-1)\operatorname{\mbox{-}type}(f\,a)$. The function $\operatorname{elim}_{V^{n}}\,P\ \phi$ is defined as follows
$\operatorname{elim}_{V^{n}}\,P\ \phi\ x:=\operatorname{tr}^{P}_{\alpha}\left(\phi\,A\,(\pi_{1}\,(\operatorname{desup}^{n}\,x))\,(\lambda\,a.\operatorname{elim}_{V^{n}}\,P\ \phi\ (f\,a,q\,a))\right)$
where $\alpha:\sup^{n}\,(\operatorname{desup}^{n}\,x)=x$ by Proposition 27. To construct the path, we use univalence to do equivalence induction. Let $Q$ be the type family
$Q:\prod_{B:\operatorname{Type}}V^{n}\simeq B\rightarrow\operatorname{Type}$
$Q\,B\,e:=\prod_{p:\prod_{b:B}P\,\left(e^{-1}\,b\right)}\prod_{b:B}\operatorname{tr}^{P}_{\alpha}\left(p\,\left(e\,\left(e^{-1}\,b\right)\right)\right)=p\,b$
where $\alpha:e^{-1}\,\left(e\,\left(e^{-1}\,b\right)\right)=e^{-1}\,b$ is the proof that $e$ is a retraction applied to the element $e^{-1}\,b$. Then, for the case $B:=V^{n}$ and $e:=\operatorname{id\mbox{-}equiv}$ we have
$Q\,V^{n}\,\operatorname{id\mbox{-}equiv}\equiv\prod_{p:\prod_{x:V^{n}}P\,(a)}\prod_{x:V^{n}}(p\,x=p\,x)$
which is inhabited by $\lambda\,p.\operatorname{refl\mbox{-}htpy}$. By univalence we thus have an element of the type $Q\,B\,e$ for any type $B$ and equivalence $e:V^{n}\simeq B$. For $B:=V^{n}$ and $e:=\operatorname{desup}^{n}$ we construct the term
$p:\prod_{(A^{\prime},f^{\prime}):\operatorname{P}^{n}_{U}V^{n}}P\,(\operatorname{sup}^{n}\,(A^{\prime},f^{\prime}))$
$p\,(A^{\prime},f^{\prime}):=\phi\,A^{\prime}\,f^{\prime}\,(\operatorname{elim}_{V^{n}}\,P\ \phi\ \circ f^{\prime})$
Then we have an element of the type $Q\,\left(\operatorname{P}^{n}_{U}V^{n}\right)\,\operatorname{desup}^{n}\,p\ (A,f)$, which unfolds to
$\left(\operatorname{tr}^{P}_{\alpha}\left(\phi\,A\,(\pi_{1}\,(\operatorname{desup}^{n}\,x))\,(\lambda\,a.\operatorname{elim}_{V^{n}}\,P\ \phi\ (f\,a,q\,a))\right)=\phi\,A\,f\,(\operatorname{elim}_{V^{n}}\,P\ \phi\ \circ f)\right)$
where $\alpha:\sup^{n}\,\left(\operatorname{desup}^{n}\,\left(\sup^{n}\,(A^{\prime},f^{\prime})\right)\right)=\sup^{n}\,(A^{\prime},f^{\prime})$ is the same element as was used in the construction of $\operatorname{elim}_{V^{n}}\,P\ \phi$.∎
As usual, we can consider the specialisation from eliminators to recursors, and in the recursors the computation rules end up holding definitionally. Since $(n-1)$-truncated maps are also functions, $V^{n}$ actually has recursors for both $\operatorname{P}^{\infty}_{U}$-algebras and $\operatorname{P}^{n}_{U}$-algebras.
Proposition 29 (Untruncated recursion for Vn ).
Given any type $X$ and map $m:\operatorname{P}^{\infty}_{U}X\rightarrow X$ there is a map $\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m:V^{n}\rightarrow X$ such that for $(A,f):\operatorname{P}^{n}_{U}V^{n}$ the following definitional equality holds:
$\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m\ (\operatorname{sup}^{n}\,(A,f))\equiv m\,(\operatorname{P}^{\infty}_{U}(\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m)\ (A,f))$
Proof.
Given $x:V^{n}$, we may assume that $x\equiv(\operatorname{sup}^{\infty}\,(A,f),(p,q))$ where $A:U$, $f:A\rightarrow V^{\infty}$, $p:\operatorname{is\mbox{-}}(n-1)\operatorname{\mbox{-}trunc\mbox{-}map\,}f$ and $q:\prod_{a:A}\operatorname{is\mbox{-}it\mbox{-}}(n-1)\operatorname{\mbox{-}type}(f\,a)$. The map $\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}$ is defined as:
$\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m\ x:=m\,(A,\lambda\,a.\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m\ (f\,a,q\,a))$
For $(A,f):\operatorname{P}^{n}_{U}V^{n}$ we thus have the following chain of definitional equalities:
$\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m\ (\operatorname{sup}^{n}\,(A,f))$ $\equiv\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m\ (\operatorname{sup}^{\infty}\,(A,\pi_{0}\circ f),(\_,\pi_{1}\circ f))$
$\equiv m\,(A,\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m\circ f)$
$\equiv m\,(\operatorname{P}^{\infty}_{U}(\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m)\ (A,f))\qed$
Corollary 5 (Truncated recursion for Vn ).
Given a $\operatorname{P}^{n}_{U}$-algebra $(X,m)$ there is a map $\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m):V^{n}\rightarrow X$ such that for $(A,f):\operatorname{P}^{n}_{U}V^{n}$ the following definitional equality holds:
$\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)\,(\operatorname{sup}^{n}\,(A,f))\equiv m\,(\operatorname{P}^{n}_{U}(\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m))\ (A,f))$
Proof.
The map $\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)$ is defined as:
$\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m):=\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ (\lambda\,(A,f).m\,(\operatorname{image}_{n-1}f,\operatorname{incl}_{n-1}f))$
For $(A,f):\operatorname{P}^{n}_{U}V^{n}$ we have the following chain of definitional equalities:
$\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)\,(\operatorname{sup}^{n}\,(A,f))$ $\equiv\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m^{\prime}\,(\operatorname{sup}^{n}\,(A,f))$
$\equiv m^{\prime}\,(A,\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m^{\prime}\circ f)$
$\equiv m\,(\operatorname{image}_{n-1}(\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m^{\prime}\circ f),\operatorname{incl}_{n-1}(\operatorname{P}^{\infty}\operatorname{\mbox{-}rec}_{V^{n}}\,X\ m^{\prime}\circ f))$
$\equiv m\,(\operatorname{image}_{n-1}(\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)\circ f),\operatorname{incl}_{n-1}(\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)\circ f))$
$\equiv m\,(\operatorname{P}^{n}_{U}(\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m))\ (A,f))$
where $m^{\prime}:=\lambda\,(A,f).m\,(\operatorname{image}_{n-1}f,\operatorname{incl}_{n-1}f)$.∎
6.3 Initiality of Vn
In the proof of initiality we are to show that the type of homomorphisms from the initial algebra, into any other algebra, is contractible. If we were working with mere sets, it would be sufficient to show that the underlying maps of the homomorphisms are equal to a specified canonical map. But being a homomorphism is actually a structure when the types involved are of higher levels. So, the proof of contractibility has to coherently transfer this structure when proving that every homomorphism is equal to the center of contraction. This is achieved by using the induction principle for $V^{n}$ and the characterisation of the identity type on $\operatorname{P}^{n}_{U}$-algebra homomorphisms.
Theorem 15 ( ).
The algebra $(V^{n},\sup^{n})$ is initial: given any other $\operatorname{P}^{n}_{U}$-algebra $(X,m)$ the type of algebra homomorphisms from $(V^{n},\sup^{n})$ to $(X,m)$ is contractible.
Proof.
The center of contraction is given by $\left(\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m),\operatorname{refl\mbox{-}htpy}\right)$. We will use the characterisation of equality between $\operatorname{P}^{n}_{U}$-algebra homomorphisms given by Lemma 8 to show that any other homomorphism is equal to $\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)$.
Let $(\phi,\alpha)$ be another $\operatorname{P}^{n}_{U}$-algebra homomorphism from $(V^{n},\sup^{n})$ to $(X,m)$. We need to do two things:
  • Construct a homotopy $\epsilon:\phi\sim\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)$.
  • For each $(A,f):\operatorname{P}^{n}_{U}\,V^{n}$, construct a path
    $\alpha\,(A,f)\cdot\operatorname{ap}_{m}\left(\operatorname{P}^{n}_{U}\epsilon\,(A,f)\right)=\epsilon\left(\operatorname{sup}^{n}\,(A,f)\right)\cdot\operatorname{refl\mbox{-}htpy}\,(A,f).$
To construct a homotopy from $\phi$ to $\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)$ we use the elimination principle on $V^{n}$, by Proposition 28, with the type family $P\,x:=\left(\phi\,x=\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)\,x\right)$. This means that given $A:U$, $f:A\hookrightarrow_{n-1}V^{n}$ and $H:\prod_{a:A}\phi\,(f\,a)=\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)\,(f\,a)$, we need to construct a path
$\phi\,(\operatorname{sup}^{n}\,(A,f))=\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)\,(\operatorname{sup}^{n}\,(A,f))$
We have the following chain of equalities:
$\phi\,(\operatorname{sup}^{n}\,(A,f))$ $=m\,(\operatorname{P}^{n}_{U}\phi\,(A,f))$ (123)
$\equiv m\,(\operatorname{image}_{n-1}(\phi\circ f),\operatorname{incl}_{n-1}(\phi\circ f))$ (124)
$=m\,(\operatorname{image}_{n-1}(\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)\circ f),\operatorname{incl}_{n-1}(\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)\circ f))$ (125)
$\equiv m\,(\operatorname{P}^{n}_{U}(\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m))\,(A,f))$ (126)
$\equiv\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)\,(\operatorname{sup}^{n}\,(A,f))$ (127)
Step (123) is the path $\alpha\,(A,f)$ and step (125) is the path $\operatorname{ap}_{\lambda\,h.m\,(\operatorname{image}_{n-1}h,\operatorname{incl}_{n-1}h)}(\operatorname{funext}\,H)$. So let
$\sigma:=\lambda\,A\,f\,H.\alpha\,(A,f)\cdot\operatorname{ap}_{\lambda\,h.m\,(\operatorname{image}_{n-1}h,\operatorname{incl}_{n-1}h)}(\operatorname{funext}\,H)$
then by Proposition 28 we have a homotopy $\operatorname{elim}_{V^{n}}\,P\ \sigma:\phi\sim\operatorname{P}^{n}\operatorname{\mbox{-}rec}_{V^{n}}\,(X,m)$.
It remains to construct the second component of the $\Sigma$-type in Lemma 8. We have the following chain of equalities:
$\alpha\,(A,f)$ $\cdot\operatorname{ap}_{m}\left(\operatorname{P}^{n}_{U}\left(\operatorname{elim}_{V^{n}}\,P\ \sigma\right)\,(A,f)\right)$ (128)
$\equiv\alpha\,(A,f)\cdot\operatorname{ap}_{m}\left(\operatorname{ap}_{\lambda\,\psi.\operatorname{P}^{n}_{U}\psi(A,f)}\left(\operatorname{funext}\,\left(\operatorname{elim}_{V^{n}}\,P\ \sigma\right)\right)\right)$ (129)
$=\alpha\,(A,f)\cdot\operatorname{ap}_{m}\left(\operatorname{ap}_{\lambda\,h.(\operatorname{image}_{n-1}h,\operatorname{incl}_{n-1}h)}\left(\operatorname{ap}_{-\circ f}\left(\operatorname{funext}\,\left(\operatorname{elim}_{V^{n}}\,P\ \sigma\right)\right)\right)\right)$ (130)
$=\alpha\,(A,f)\cdot\operatorname{ap}_{\lambda\,h.m\,(\operatorname{image}_{n-1}h,\operatorname{incl}_{n-1}h)}\left(\operatorname{ap}_{-\circ f}\left(\operatorname{funext}\,\left(\operatorname{elim}_{V^{n}}\,P\ \sigma\right)\right)\right)$ (131)
$=\alpha\,(A,f)\cdot\operatorname{ap}_{\lambda\,h.m\,(\operatorname{image}_{n-1}h,\operatorname{incl}_{n-1}h)}\left(\operatorname{funext}\,\left(\left(\operatorname{elim}_{V^{n}}\,P\ \sigma\right)\circ f\right)\right)$ (132)
$\equiv\sigma\,A\,f\,\left(\left(\operatorname{elim}_{V^{n}}\,P\ \sigma\right)\circ f\right)$ (133)
$=\operatorname{elim}_{V^{n}}\,P\ \sigma\left(\operatorname{sup}^{n}\,(A,f)\right)$ (134)
$=\operatorname{elim}_{V^{n}}\,P\ \sigma\left(\operatorname{sup}^{n}\,(A,f)\right)\cdot\operatorname{refl\mbox{-}htpy}\,(A,f)$ (135)
In steps (130) and (131) we use the fact that $\operatorname{ap}$ and function composition commute. Step (132) uses the fact that function extensionality respects precomposition. Finally, step (134) is the computation rule for $\operatorname{elim}_{V^{n}}$.∎
6.4 Properties
Since $V^{n}$ is a fixed-point for $\operatorname{P}^{n}_{U}$, the induced $\in$-structure satisfies the properties as shown in Section 5. But as $V^{n}$ is the initial algebra for $\operatorname{P}^{n}_{U}$, it also satisfies foundation, which is not true of all fixed-points for $\operatorname{P}^{n}_{U}$. More explicitly, we list the properties which $(V^{n},\in^{n})$ satisfies.
Theorem 16 ( ).
For $n:\mathbb{N}^{\infty}$, the $\in$-structure $(V^{n},\in^{n})$ satisfies the following properties:
  • empty set,
  • $U$-restricted $n$-separation,
  • $\infty$-unordered $I$-tupling, for all $k:\mathbb{N}_{-1}$ such that $k<n$ and $k$-truncated types $I:U$,
  • $k$-unordered $I$-tupling, for all $k:\mathbb{N}_{-1}$ such that $k\leq n$ and $I:U$,
  • $k$-replacement, for all $k:\mathbb{N}_{-1}$ such that $k\leq n$,
  • $k$-union, for all $k:\mathbb{N}_{-1}$ such that $k\leq n$,
  • exponentiation, for any ordered pairing structure,
  • natural numbers for any $(n-1)$-truncated representation,
  • foundation.
Proof.
The first two properties follow directly from Theorem 4 and Theorem 5 respectively since $V^{n}$ is a fixed-point for $\operatorname{P}^{n}_{U}$. The next four properties follow from Theorem 8, Theorem 6, Theorem 9 and Theorem 10 respectively, together with the fact that $V^{n}$ is locally small (Proposition 24), and thus $(k+1)$-small for all $k:\mathbb{N}$.
By Theorem 11, $(V^{n},\in^{n})$ has exponentiation for any ordered pairing structure. (Note that $(V^{n},\in^{n})$ has at least one ordered pairing structure by Theorem 7.)
For natural numbers, the result follows from Theorem 12, since $V^{n}$ is a fixed-point for $\operatorname{P}^{n}_{U}$.
Lastly, we use the induction principle for $V^{n}$ to show that $(V^{n},\in^{n})$ has foundation. Given $A:U$ and $f:A\hookrightarrow_{n-1}V^{n}$ we need to construct an element of type
$\left(\prod_{a:A}\operatorname{Acc}\,(f\,a)\right)\rightarrow\operatorname{Acc}\,(\operatorname{sup}^{n}\,(A,f)).$
Therefore suppose we have $p:\prod_{a:A}\operatorname{Acc}\,(f\,a)$. We construct the following element:
$\operatorname{acc}\,\left(\lambda\,y\,(a,q).\operatorname{tr}^{Acc}_{q}(p\,a)\right):\operatorname{Acc}\,(\operatorname{sup}^{n}\,(A,f)).$
It then follows by Proposition 28 that we have
$\prod_{x:V^{n}}\operatorname{Acc}\,x.\qed$
7 $\operatorname{V}^{n}$ as an $n$-type universe of $n$-types
[…]
Let $(\operatorname{V}^{n},\in^{n})$ be the $n$-level $\in$-structure given by Theorem 14. Note that the type $x\in^{n}y$ is an $(n-1)$-type, for any elements $x,y:\operatorname{V}^{n}$.
Proposition 30 ( ).
The type $\operatorname{V}^{n}$ is an $n$-type.
Proof.
This follows both from Proposition 1, and from Proposition 23 together with Theorem 13.∎
Definition 31 ( ).
For $n:\mathbb{N}^{\infty}$, define the decoding function on $\operatorname{V}^{n}$ as the family
$\operatorname{El}^{n}:\operatorname{V}^{n}\rightarrow U$
$\operatorname{El}^{n}\,x:=\overline{x}$
We take this as the decoding function rather than $\operatorname{El}$ from Definition 2. This is so that the decoding holds up to definitional equality, since we have
$\operatorname{El}^{n}\,(\operatorname{sup}^{n}\,(A,f))\equiv A.$
But the two families are equivalent by Proposition 22.
Proposition 31 ( ).
For $n:\mathbb{N}^{\infty}$, the decoding $\operatorname{El}^{n}\,a$ of any element $a:\operatorname{V}^{n}$, is an $n$-type.
Proof.
The map $\widetilde{a}$ is an $(n-1)$-truncated map from $\operatorname{El}^{n}\,a$ into $\operatorname{V}^{n}$. The domain $\operatorname{El}^{n}\,a$ is equivalent to the total space of fibers of $\widetilde{a}$, which is an $n$-type as the base, $\operatorname{V}^{n}$, is an $n$-type and each fiber is an $(n-1)$-type.∎
Propositions 30 and 31 thus show that $\operatorname{V}^{n}$ is an $n$-type universe of $n$-types. Section 3 explored internalisations of types in $\in$-structures. The type of internalisations of a type was shown to be equivalent to the type of internalisable representations of that type (Proposition 11). In the case of $\operatorname{V}^{n}$, the internalisable representations are the $(n-1)$-truncated ones.
Proposition 32 ( ).
For $n:\mathbb{N}^{\infty}$ and a type $A:U$ together with a representation $f:A\rightarrow\operatorname{V}^{n}$, the representation is internalisable if and only if it is $(n-1)$-truncated.
Proof.
Suppose that $f$ is internalisable, i.e. we have an element of the type
$\sum_{a:\operatorname{V}^{n}}\prod_{z:\operatorname{V}^{n}}z\in^{n}a\simeq\operatorname{fiber}f\,z.$
Since the type $z\in^{n}a$ is an $(n-1)$-type for all $z:\operatorname{V}^{n}$ it follows that $\operatorname{fiber}f\,z$ is an $(n-1)$-type for all $z:\operatorname{V}^{n}$. Conversely, suppose that $f$ is $(n-1)$-truncated. Then we take for $a$ the element $\sup^{n}\,(A,f)$ and for the family of equivalences, the family of identity equivalences.∎
Proposition 33 ( ).
For $n:\mathbb{N}^{\infty}$ and $A:U$, the type of $(n-1)$-truncated representations of $A$, $A\hookrightarrow_{n-1}\operatorname{V}^{n}$, is equivalent to the type of internalisations of $A$, $\sum_{a:A}\operatorname{El}^{n}\,a\simeq A$.
Proof.
This follows by Proposition 12 and the fact that the families $\operatorname{El}$ and $\operatorname{El}^{n}$ are equivalent, together with the previous proposition. Alternatively, $\sup^{n}$ is an equivalence making the following diagram commute
[…]
Therefore the type of fibers of $\pi_{0}$ over $A$, which is $A\hookrightarrow_{n-1}\operatorname{V}^{n}$, is equivalent to the type of fibers of $\operatorname{El}^{n}$ over $A$, which is equivalent to $\sum_{a:A}\operatorname{El}^{n}\,a\simeq A$ by univalence.∎
The universes are cumulative both with regards to universe levels and with regards to type levels. So far we have assumed only two universes, $U$ and $\operatorname{Type}$, but assume for the next proposition a (cumulative) hierarchy of universes $U_{0},U_{1},\cdots,U_{\ell},\cdots$ and let $\operatorname{V}^{n}_{\ell}$ denote the initial algebra to the functor $\operatorname{P}^{n}_{U_{\ell}}$.
Proposition 34 ( ).
For $n:\mathbb{N}^{\infty}$, there is an internalisation of $\operatorname{V}^{n}_{\ell}$ in $\operatorname{V}^{n}_{\ell+1}$.
Proof.
The map
$\phi:\operatorname{V}^{\infty}_{\ell}\rightarrow\operatorname{V}^{\infty}_{\ell+1}$
$\phi\,(\operatorname{sup}^{\infty}\,(A,f)):=\operatorname{sup}^{\infty}\,(A,\phi\circ f)$
[…]
The map $\phi$ is therefore an embedding $\operatorname{V}^{n}_{\ell}\hookrightarrow\operatorname{V}^{n}_{\ell+1}$. By Proposition 33 the map gives rise to an internalisation of $\operatorname{V}^{n}_{\ell}$, namely $\sup^{n+1}\,(\operatorname{V}^{n}_{\ell},\phi)$.∎
Proposition 35 ( ).
For $n:\mathbb{N}^{\infty}$, there is an embedding $\operatorname{V}^{n}\hookrightarrow\operatorname{V}^{n+1}$.
Proof.
This is simply the fact that an iterative $n$-type is also an iterative $(n+1)$-type.∎
Note that there is a size issue with internalising $\operatorname{V}^{n}$ in $\operatorname{V}^{n+1}$. If there was an element $v:\operatorname{V}^{n+1}$ such that $\operatorname{El}^{n+1}\,v\simeq\operatorname{V}^{n}$, then the type $\operatorname{V}^{n}$ would be essentially $U$-small, which would induce a paradox. However, assuming again a hierarchy of universes, the type $\operatorname{V}^{n}_{\ell}$ can be internalised in $\operatorname{V}^{n+1}_{\ell+1}$.
The universe $\operatorname{V}^{n}$ also contains all the usual types and type formers, assuming they exist in the underlying universe $U$.
Proposition 36 ( ).
For $n:\mathbb{N}^{\infty}$, the universe $\operatorname{V}^{n}$ contains the following types and type formers:
  • the empty type, unit type and booleans,
  • the natural numbers,
  • $\Pi$-types,
  • $\Sigma$-types,
  • coproducts and
  • identity types.
Proof.
[…]
For $\Pi$-types and $\Sigma$-types, suppose we have an element $a:\operatorname{V}^{n}$ and a map $b:\operatorname{El}^{n}\,a\rightarrow\operatorname{V}^{n}$. In the first case, note that $\operatorname{graph}_{\,\widetilde{a},\lambda\,i.\widetilde{(b\,i)}}$ in Lemma 12 gives an $(n-1)$-truncated representation of $\prod_{i:\operatorname{El}^{n}\,a}\operatorname{El}^{n}\,(b\,i)$. For $\Sigma$-types, we use the ordered pairing structure given by Theorem 7. The map $\lambda\,(i,j).\langle\widetilde{a}\,i,\widetilde{(b\,i)}\,j\rangle$ is an $(n-1)$-truncated representation of $\sum_{i:\operatorname{El}^{n}\,a}\operatorname{El}^{n}\,(b\,i)$.
Given $a,b:\operatorname{V}^{n}$, the map
$f:\operatorname{El}^{n}\,a+\operatorname{El}^{n}\,b\rightarrow\operatorname{V}^{n}$
$f\,(\operatorname{inl}\,i):=\langle\emptyset,\widetilde{a}\,i\rangle$
$f\,(\operatorname{inr}\,j):=\langle\{\emptyset\}_{0},\widetilde{b}\,j\rangle$
is $(n-1)$-truncated. To see this, note that for $z:\operatorname{V}^{n}$ we have the following chain of equivalences:
$\operatorname{fiber}\,f\,z$ $\equiv\left(\sum_{s:\operatorname{El}^{n}\,a+\operatorname{El}^{n}\,b}f\,s=z\right)$ (136)
$\simeq\left(\sum_{i:\operatorname{El}^{n}\,a}\langle\emptyset,\widetilde{a}\,i\rangle=z\right)+\left(\sum_{j:\operatorname{El}^{n}\,b}\langle\{\emptyset\}_{0},\widetilde{b}\,j\rangle=z\right)$ (137)
$\simeq\left(\sum_{((s,t),p):\operatorname{fiber}\,\langle-,-\rangle\,z}\left(\operatorname{fiber}\,\widetilde{a}\,t\right)\times\left(s=\emptyset\right)\right)$ (138)
$\hskip 10.0pt+\left(\sum_{((s,t),p):\operatorname{fiber}\,\langle-,-\rangle\,z}\left(\operatorname{fiber}\,\widetilde{b}\,t\right)\times\left(s=\{\emptyset\}_{0}\right)\right)$
The last type is $(n-1)$-truncated as all the components are $(n-1)$-truncated and the two summands are mutually exclusive since {}$\emptyset\neq\{\emptyset\}_{0}$. The map $f$ therefore gives an internalisation of the coproduct $\operatorname{El}^{n}\,a+\operatorname{El}^{n}\,b$. For identity types we note that the map $\lambda\,p.\,\emptyset:a=b\rightarrow\operatorname{V}^{n}$ is $(n-1)$-truncated as it is a function from an $(n-1)$-type into an $n$-type. So the type $a=b$ is internalisable.∎
Note that all the codes in $\operatorname{V}^{n}$ are constructed using $\sup^{n}$. Therefore, the decoding of a code is definitionally equal to the type being encoded.
As $\operatorname{V}^{n}$ is an $n$-type universe of $n$-types, a natural question to ask is which $n$-types lie in the universe. For $\operatorname{V}^{1}$ we have seen that the classifying space of any group can be internalised. In particular, we have an internalisation of the higher inductive type of the circle in $\operatorname{V}^{1}$. Using the same idea, we can internalise any $n$-type in $\operatorname{V}^{n+1}$.
Proposition 37 ( ).
For $n:\mathbb{N}^{\infty}$, any $U$-small $n$-type can be internalised in $\operatorname{V}^{n+1}$.
Proof.
Let $A:U$ be an $n$-type. Let $f:A\rightarrow\operatorname{V}^{n+1}$ be the constant function sending any element to $\emptyset$. This is an $n$-truncated map as the domain is an $n$-type and the codomain is an $(n+1)$-type. The element $\sup^{n+1}\,(A,f)$ is thus an internalisation of $A$ in $\operatorname{V}^{n+1}$.∎
For $\operatorname{V}^{0}$, the statement that any mere set has an internalisation is essentially the axiom of wellfounded materialisation [shulman_stack_2010]. The statement that any $n$-type can be internalised in $\operatorname{V}^{n}$ can thus be thought of as a higher level generalisation of wellfounded materialisation. It amounts to being able to equip every $n$-type with some higher iterative structure.
Proposition 38.
For $n:\mathbb{N}^{\infty}$, the universe $\operatorname{V}^{n}$ is not a univalent universe.
Proof.
The quickest way to see this is to note that $\operatorname{V}^{n}$ contains (at least) two distinct internalisations of the unit type: {}$\{\emptyset\}_{0}$ and {{}$\{\{\emptyset\}_{0}\}_{0}$ (but there are of course many more). As these both decode to the unit type but are not equal, univalence fails.∎
[…]
8 Conclusion and future work
In this paper we defined the concept of an $\in$-structure and gave generalisations of the axioms of constructive set theory to higher level structures. As instances of such higher level $\in$-structures, we generalised the construction of the type of iterative sets to obtain the initial algebras, $\operatorname{V}^{n}$, to the functors $\operatorname{P}^{n}_{U}$. These were shown to model the $\in$-structure properties at the same level or lower. Moreover, $\operatorname{V}^{n}$ was shown to be an $n$-type universe of $n$-types with definitional decoding.
8.1 Related work
Gallozzi [gallozzi] constructs a family of interpretations of set theory in homotopy type theory indexed on two type levels, $k$ and $h$. His interpretation is a generalisation of Aczel’s model. In one direction, he generalises by taking as the type of sets Aczel’s W-type, but over the small universe of $k$-types, rather than the whole universe $U$. In the other direction, he uses $h$-truncated $\Sigma$-types, while Aczel uses untruncated $\Sigma$-types. Gallozzi then shows that this models Aczel’s CZF [aczel1978] and Myhill’s CST [myhill_constructive_1975].
These models are setoid models — equality is not interpreted as the identity type, as opposed to our models. They are not $\in$-structures in our sense either, in that our version of extensionality does not hold. Of course, he shows that his interpretation of extensionality holds. The $k$-level W-type used by Gallozzi is a $(k+1)$-type, and it is a subtype of our type $\operatorname{V}^{k+1}$, so any set in that universe is also a set in our universe.
The HoTT Book model of set theory is equivalent to the iterative sets model. It is not clear, however, how to generalise that construction as we have done with the iterative sets, as this would require more complex higher inductive types.
8.2 Future work
As we have seen, the universe $\operatorname{V}^{n}$ embeds into the universe $\operatorname{V}^{n+1}$. So the higher universes contain the types of the lower universes. However, it remains to see which new types appear at each level. For instance, we showed that the higher inductive type of the circle, and more generally the classifying space of any group, lies in $\operatorname{V}^{1}$. For the universe $\operatorname{V}^{n}$ one would like to construct an internalisation of (at least some) proper $n$-types. This amounts to giving the type some kind of higher iterative structure, which would be an interesting direction for further study.
[…]